### v1

# Knapsack Optimization-based Schema Linking for LLM-based Text-to-SQL Generation

Abstract
Generating SQLs from user queries is a long-standing challenge, where the accuracy of initial schema linking significantly impacts subsequent SQL generation performance. However, current schema linking models still struggle with missing relevant schema elements or an excess of redundant ones. A crucial reason for this is that commonly used metrics, recall and precision, fail to capture relevant element missing and thus cannot reflect actual schema linking performance. Motivated by this, we propose an enhanced schema linking metric by introducing a restricted missing indicator. Accordingly, we introduce Knapsack optimization-based Schema Linking Agent (KaSLA), a plug-in schema linking agent designed to prevent the missing of relevant schema elements while minimizing the inclusion of redundant ones. KaSLA employs a hierarchical linking strategy that first identifies the optimal table linking and subsequently links columns within the selected table to reduce linking candidate space. In each linking process, it utilize a knapsack optimization approach to link potentially relevant elements while accounting for a limited tolerance of potential redundant ones. With this optimization, KaSLA-1.6B achieves superior schema linking results compared to large-scale LLMs, including deepseek-v3 with state-of-the-art (SOTA) schema linking method. Extensive experiments on Spider and BIRD benchmarks verify that KaSLA can significantly improve the SQL generation performance of SOTA text-to-SQL models by substituting their schema linking processes.
Knapsack Optimization-based Schema Linking for
LLM-based Text-to-SQL Generation
Zheng Yuan1, Hao Chen2, Zijin Hong1, Qinggang Zhang1 Feiran Huang3, Xiao Huang1 1The Hong Kong Polytechnic University 2City University of Macau, 3Jinan University {yzheng.yuan, zijin.hong, qinggangg.zhang}@connect.polyu.hk sundaychenhao@gmail.com; huangfr@jnu.edu.cn; xiao.huang@polyu.edu.hk
1 Introduction
Figure 1: In text-to-SQL tasks, commonly used schema linking metrics such as Recall and Precision fail to accurately capture the true performance of schema linking when relevant schema elements are missing.
Figure 2: Overall framework of proposed knapsack optimization-based schema linking agent (KaSLA) with column linking as an example. Traditional schema linking models often struggle with both missing relevant elements and the inclusion of redundant ones. KaSLA addresses this challenge by maximizing the total estimated relevance while maintaining total redundancy within a specified tolerance using knapsack optimization. Serving as a general model, KaSLA enhances the SQL generation of text-to-SQL models by replacing their existing schema linking processes.
With the advent of large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024), LLM-based text-to-SQL is emerging as the next-generation interface for database users (Hong et al., 2024b). Typically, text-to-SQL frameworks employ a two-step process: first linking database schema elements (tables and columns) to user queries, then generating the corresponding SQL statements. However, the accurate SQL generation requires the linking to the correct schema elements, meaning that errors in early stages are inevitably propagated to later SQL generation. For instance, inaccurate schema linking—such as overlooking relevant schema elements or linking to irrelevant ones—will lead to incorrect SQL statements.
Recent state-of-the-art text-to-SQL models tend to focus primarily on final SQL generation while employing relatively simplistic schema linking strategies or neglecting it. DTS-SQL (Pourreza and Rafiei, 2024) finetunes a local LLM to select relevant tables based on user queries. Solid-SQL (Liu et al., 2025) present a data augmentation method for the finetuning of Schema linking method. Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) use semantic-based retrieval models to enhance the schema linking process by retrieving relevant information from databases. RSL-SQL (Cao et al., 2024) introduces a bidirectional approach to simplify the full schema. E-SQL (Qu et al., 2024) and (Maamari et al., 2024) assert that schema linking remains challenging for current LLMs and thus directly use the full schema as input for SQL generation.
Even though schema linking is crucial, current widely used metrics, such as Precision and Recall (Li et al., 2023a, 2024b; Qu et al., 2024), cannot adequately evaluate instances when relevant elements are missing. As demonstrated in Figure 1, missing a single relevant table or column can significantly decrease the accuracy of SQL generation, yet still result in high Recall and Precision scores. Because Recall does not significantly penalize the model as long as the majority of the relevant elements are identified, allowing it to achieve an artificially high score, even when crucial elements are missing. Similarly, regardless of the severity of the missing, as long as the number of correctly linked elements is proportionally high compared to the total predicted relevant elements, the model can achieve an unreasonably high Precision score. This discrepancy distorts the assessment of a schema linking model’s effectiveness, potentially hindering progress in text-to-SQL development.
To address this, we firstly design a tailored schema linking metrics that accurately assess linking performance with a restricted element missing indicator. Additionally, we propose the Knapsack optimization-based Schema Linking Agent (KaSLA), designed to link the most relevant and minimally redundant schema elements according to the given queries. KaSLA introduces a binary-probabilistic score function to predict a robust relevance score of each element. Additionally, KaSLA identifies redundancy of each element and the upper redundancy tolerance for each query based on its most similar queries in the training dataset, effectively minimizing the inclusion of redundant elements. Finally, KaSLA links tables and columns with an hierarchical strategy to reduce the column linking candidate space for efficiency and employs a 0-1 knapsack optimization approach to select the most relevant element set under the upper redundancy tolerance to reduce the redundant linking. Extensive experiments demonstrate that KaSLA can improve SQL generation accuracy by reducing missing and redundant information in the schema linking.
In summary, our contributions are as follows:
  • We introduce an enhanced schema linking metric, which can accurately capture the missing of relevant elements and reflect the actual linking performance aligning with the SQL generation performance.
  • We propose the Knapsack optimization-based Schema Linking Agent (KaSLA) to ensure the linking of relevant schema elements and minimizing the inclusion of redundant ones.
  • KaSLA can serve as a general agent to enhance LLM-based text-to-SQL framework by provide optimal linking results. Extensive experiments demonstrate improved performance in the SOTA text-to-SQL models when using KaSLA for schema linking.
2 Preliminaries
In this section, we outline the definitions for the schema linking and 0-1 Knapsack Optimization.
Schema Linking Task.
Given a natural language query $q$ and a corresponding database schema ${\mathcal{S}}_{q}$ containing schema elements, tables ${\mathcal{T}}=\{t_{1},\cdots,t_{|{\mathcal{T}}|}\}$ and columns ${\mathcal{C}}=\{c_{1},\cdots,c_{|{\mathcal{C}}|}\}$ in each table. Schema linking is a common process used in text-to-SQL to predict the relevant schema entities with $q$ from ${\mathcal{S}}_{q}$:
$f_{\text{linking}}:(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q},$ (1)
where $\widehat{{\bm{S}}}_{q}$ is the subset of ${\mathcal{S}}_{q}$ that includes schema element $s\in{\mathcal{S}}_{q}$ that relevant to $q$. Each $s$ could represent a specific table or column. Additionally, we denote ${\bm{S}}_{q}^{*}$ as the ground truth linking results.
0-1 Knapsack Optimization.
The 0-1 knapsack problem Fréville (2004) is a classic combinatorial optimization challenge. It involves selecting items from a finite item set, each with a value ${\mathcal{V}}$ and weight ${\mathcal{W}}$, to place into a knapsack with a maximum weight capacity ${\mathcal{C}}$. Each item can be selected only once. Formally, the objective of 0-1 knapsack optimization is to select items such that their total weight remains below ${\mathcal{C}}$ while maximizing total value of these selected items:
$\widehat{N}=\arg\max\sum_{i\in\widehat{N}}{\mathcal{V}}_{i},\ \text{s.t.}\sum_{i\in\widehat{N}}{\mathcal{W}}_{i}\leq{\mathcal{C}},$ (2)
where $\widehat{N}$ is the selected item set for the knapsack.
3 Enhanced Schema Linking Metrics
In this section, we introduce enhanced schema linking metrics considering the missing of relevant elements during evaluation, thereby accurately assess the true performance of schema linking models.
For a query $q$ with predicted schema linking result $\widehat{{\bm{S}}}_{q}$ and ground truth linking result ${\bm{S}}^{*}_{q}$, Recall and Precision are two commonly used metreics (Li et al., 2023a, 2024b; Qu et al., 2024). Recall, ${\bm{R}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|{\bm{S}}^{*}_{q}|}$, measures the proportion of relevant elements correctly identified, specifically how many of the actual relevant elements in ${\bm{S}}^{*}_{q}$ are found in $\widehat{{\bm{S}}}_{q}$. Conversely, Precision, ${\bm{P}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|\widehat{{\bm{S}}}_{q}|}$, assesses the accuracy of predictions by determining the proportion of elements predicted as relevant in $\widehat{{\bm{S}}}_{q}$ that are indeed correct.
As illustrated in Figure 1, these metrics can yield high scores even in the missing of relevant elements. This underscores that they do not effectively address the issue of missing relevant elements, which can undermine the true accuracy of schema linking models.
To address the limitations of Recall and Precision, we firstly introduce a restricted element missing indicator to detect whether relevant elements are missing in the predicted schema linking result $\widehat{{\bm{S}}}_{q}$, The formal definition is as follows:
$\bm{\mathbbm{1}}(q)=\begin{cases}1,&\text{if }\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}={\bm{S}}^{*}_{q},\\ 0,&\text{otherwise}.\end{cases}$ (3)
It takes a value of 1 when the predicted result $\widehat{{\bm{S}}}_{q}$ contain all the elements in the ground truth ${\bm{S}}^{*}_{q}$, meaning no relevant elements are missing. Conversely, it is set to 0 when there are any missing. Building on this missing indicator, we define enhanced Recall and Precision as follows:
${\bm{R}}^{\text{{+}}}(q)=\bm{\mathbbm{1}}(q){\bm{R}}(q),{\bm{P}}^{\text{{+}}}(q)=\bm{\mathbbm{1}}(q){\bm{P}}(q).$ (4)
Such metrics strengthen the penalty for missing relevant elements, as schema linking results with missing are inherently incorrect for SQL generation. For simplicity, we use the harmonic mean of enhanced Recall and Precision, referred to as the enhanced F1 score, as the overall measure for evaluating schema linking results:
${{\bm{F}}1}^{\text{{+}}}(q)=\frac{2{\bm{R}}^{\text{{+}}}(q){\bm{P}}^{\text{{+}}}(q)}{{\bm{R}}^{\text{{+}}}(q)+{\bm{P}}^{\text{{+}}}(q)}.$ (5)
This approach provides a comprehensive evaluation of both recall and precision while incorporating the critical aspect of missing elements.
4 Knapsack Optimization-based Schema Linking Agent (KaSLA)
In this section, we introduce the proposed KaSLA framework, which focuses on ensuring the linking of relevant schema elements to avoid missing and minimizing the inclusion of redundant ones. As illustrated in Figure 2, We begin by detailing a binary-probabilistic score function used for assessing the relevance of schema elements and an estimation function for the redundancy tolerance of each schema linking process. We then outline the knapsack optimization approach to link potential relevant element while reducing the potential redundant ones under the tolerance constraint. Finally, we describe the hierarchical schema linking process, demonstrating its efficiency for reducing the column linking candidate.
4.1 Relevance Estimation
To precisely assess the relevance of each schema element $s$ in relation to the given query $q$, we employ a hybrid binary-probabilistic score function. This function integrates a generative model for binary scoring alongside an encoding model for probabilistic scoring, allowing for a comprehensive estimation of relevance.
Binary scoring function.
This component is tasked with providing binary relevance assessments, determining whether a schema element is relevant or not in a straightforward manner:
$f_{\text{binary}}:(q,{\mathcal{S}}_{q})\mapsto\{r^{b}_{1},\cdots,r^{b}_{|{\mathcal{S}}_{q}|}\},$ (6)
where $r^{b}_{i}\in\{0,1\}$ represents the binary score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$. A score of 1 indicates relevant while 0 signifies redundant.
Considering efficiency, we finetune a lightweight LLM, DeepSeek-Coder-1.3B (Guo et al., 2024), using LoRA (Hu et al., 2021), a well-established parameter-efficient technique. We employ a strategy whereby the initial generation of a simulated SQL provides a contextual foundation for the subsequent generation of schema linking. The loss function during fine-tuning is defined as follows:
$\mathcal{L}_{\text{binary}}=-\sum_{q\in Q}\log P({\text{SQL}}^{*}_{q},{\bm{S}}_{q}^{*}\mid q,{\mathcal{S}}_{q}),$ (7)
where $Q$ is the set of queries in the training dataset. ${\text{SQL}}^{*}_{q}$ and ${\bm{S}}^{*}$ are the ground truth SQL and linking result, respectively. This formulation utilizes a joint generation of a simulated SQL and the schema linking result, thereby ensuring that the linking generation benefits from the contextual alignment provided by the simulated SQL during both fine-tuning and inference. Noted that the simulated SQL generated during this process will not be utilized in subsequent steps. The input and output examples are provided in Table 6 in Appendix B.
Probabilistic scoring function.
In conjunction with the binary scoring model, we introduce a probabilistic scoring model to enhance prediction robustness. The primary motivation is to mitigate the risk of the binary model misclassifing relevant elements. The probabilistic scoring model assigns a soft score to each element as the linking probability to reduce the likelihood of relevant elements missing , thereby supporting the binary model:
$f_{\text{prob.}}:(q,{\mathcal{S}}_{q})\mapsto\{r^{p}_{1},\cdots,r^{p}_{|{\mathcal{S}}_{q}|}\},$ (8)
where $r^{p}_{i}\in[0,1]$ represents the probabilistic score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$.
For the probabilistic scoring model, we utilize an encoding model, RoBERTa-Large (Liu et al., 2019), to derive the semantic embeddings for the query and all schema elements. Following the methodology of (Li et al., 2023a, 2024b), we employ a cross-attention network to jointly embed the semantic representations of tables and their columns. The dot product of the query embedding and the element embedding yields the final probabilistic score. This probabilistic model is trained using the presence of $s$ in ${\bm{S}}^{*}$ as the learning objective. The loss function is defined as follows:
$\mathcal{L}_{\text{prob.}}=\sum_{q\in Q}\text{FL}(\mathbbm{1}(s\in{\bm{S}}^{*}_{q}),f_{\text{prob.}}(q,{\mathcal{S}}_{q})),$ (9)
where $\text{FL}(\cdot)$ denotes the focal loss function (Ross and Dollár, 2017), designed to emphasize learning from hard negative samples.
Relevance scoring estimation.
The final function for estimating the relevance of each schema element integrates the scores derived from both the binary and probabilistic functions:
$\begin{split}f_{\text{relevance}}:&(q,{\mathcal{S}}_{q})\mapsto\{r_{1},\ldots,r_{|{\mathcal{S}}_{q}|}\}\\ &=\left\{\min(1,r^{b}_{i}+r^{p}_{i})\mid s_{i}\in{\mathcal{S}}_{q}\right\},\end{split}$ (10)
where $r_{i}$ denotes the relevance score assigned to the $i$-th schema element $s_{i}\in{\mathcal{S}}_{q}$ given the query $q$. An upper limit of 1 is imposed on $r_{i}$ to prevent excessively strong contrasts between potentially relevant elements, as each relevant schema element is equally important for accurate SQL generation.
4.2 Redundancy Estimation
To minimize the inclusion of redundant elements during schema linking, we define both a redundancy score for each schema element $s\in{\mathcal{S}}_{q}$ and the upper redundancy tolerance for query $q$.
Redundancy scoring estimation.
Recognizing that elements with lower relevance scores are more likely to be redundant, whereas those with higher scores are more essential, we assign a redundancy score $w_{i}$ to each schema element $s_{i}\in{\mathcal{S}}_{q}$ in relation to query $q$. This score is calculated as the inverse of its relevance score $r_{i}$, formulated as follows:
$w_{i}={r_{i}}^{-1}.$ (11)
It ensures elements with lower relevance scores contribute more to redundancy, thereby effectively prioritizing more relevant schema elements.
Upper Redundancy Tolerance.
Redundancy tolerance defines the limit on the total redundancy score of elements in the schema linking results, as an excess of redundant elements can significantly disrupt subsequent SQL generation. Estimating it during inference is challenging. For the SQL generation of some queries, a lenient tolerance may be beneficial since it allowing the inclusion of more potentially relevant elements in the linking results to prevent missing. However, for other queries, a lenient tolerance could introduce to many redundant elements, which may significantly disrupt the final SQL generation process.
To address this challenge, we hypothesize that similar queries exhibit comparable tolerance levels for redundancy and leverage queries from the training dataset, where ground truth linking results are available, to aid in predicting redundancy tolerance for given query $q$ during inference. We first retrieve the top-$K$ queries in the training dataset that closely resemble $q$ using sentence similarity measures (Gao et al., 2021). For a query $q_{k}$ in the retrieved top-$K$ queries, we calculate the total redundancy score for its ground truth linking results ${\mathcal{S}}^{*}_{q_{k}}$ and use it as its upper redundancy tolerance. The maximum of these upper redundancy tolerances from the $K$ most similar queries in the training dataset is then used as the redundancy tolerance estimate for query $q$:
$u_{q}=\max_{q_{k}\in\text{TopK}(q)}\bigg{(}\sum_{s_{j}\in{\mathcal{S}}^{*}_{q_{k}}}w_{j}\bigg{)}$ (12)
By capturing the maximum sum from the most similar queries, this approach effectively defines an upper bound on redundancy tolerance that is both efficient and adaptable, ensuring that while redundant elements are excluded, the risk of missing relevant elements is minimized.
4.3 Schema Linking with Knapsack Optimization
Given the relevance score $r_{i}$, the redundancy score $w_{i}$ for each $s_{i}\in{\mathcal{S}}_{q}$, and the upper redundancy tolerance $u_{q}$ for query $q$, we introduce a hierarchical schema linking strategy with knapsack optimization. This approach is designed to ensure the linking of relevant schema elements to avoid missing and minimize the inclusion of redundant ones for a precise schema linking.
Knapsack Optimization
The 0-1 knapsack problem (Fréville, 2004) is a classic combinatorial optimization challenge. It involves selecting items from a finite set, each item characterized by a specific value and weight, to place into a knapsack with a limited weight capacity. The objective of 0-1 knapsack optimization is to choose items that maximize the total value while ensuring the total weight remains below the knapsack’s capacity.
Inspired by this, we develop a knapsack optimization-based schema linking strategy to ensure the inclusion of relevant schema elements while minimizing redundant ones, thus ach ieving optimal schema linking. Formally, the optimization objective are defined as follows:
$\begin{split}f_{\text{Knap}}:&(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q}=\arg\max\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}r_{i}\\ &\text{s.t.}\quad\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}w_{i}\leq u_{q}.\end{split}$ (13)
This formulation aims to maximize the total relevance scores of the linked schema elements. The constraint ensures that the cumulative redundancy score remains within the upper redundancy tolerance of query $q$. To tackle this optimization problem and achieve the optimal linking result, We employ an efficient dynamic programming approach. Detailed pseudocode of this dynamic programming algorithm is provided in Algorithm A in Appendix.
Hierarchical Linking Strategy
For efficiency, our KaSLA employs a hierarchical process in schema linking: initially linking tables and subsequently linking columns within each selected table. This strategy effectively reduces the dimensionality of candidate columns, enhancing KaSLA’s capacity to handle large-scale schema linking and text-to-SQL generation tasks.
In the initial phase for a given query $q$, KaSLA aims to identify the relevant tables from the set of all available tables in schema ${\mathcal{S}}_{q}$, denoted as ${\mathcal{T}}_{q}$:
$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}(q,{\mathcal{T}}_{q}).$ (14)
After determining tables, for each $t\in\widehat{{\bm{T}}}_{q}$, let ${\mathcal{C}}^{t}_{q}$ represent the columns in table $t$. KaSLA then optimizes column linking within each selected table:
$\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}(q,{\mathcal{C}}^{t}_{q}).$ (15)
The final schema linking result $\widehat{{\bm{S}}}_{q}$ is the union of all selected tables and their corresponding selected columns, represented as:
$\widehat{{\bm{S}}}_{q}=\widehat{{\bm{T}}}_{q}\cup\bigcup_{t\in\widehat{{\bm{T}}}_{q}}\widehat{{\bm{C}}}^{t}_{q}.$ (16)
We provide a complexity analysis is detailed in Appendix D, offering insights into KaSLA’s implementation and computational efficiency.
5 Experiments
We conducted comprehensive experiments to evaluate KaSLA and address the following research questions: RQ1: Does KaSLA enhance existing text-to-SQL baselines by improving their schema linking capabilities? RQ2: Does KaSLA demonstrate superior schema linking performance by handling element missing and redundancy? RQ3: Does each component in KaSLA contribute to the overall performance, and is KaSLA an efficient model for real-world text-to-SQL applications?
5.1 Experiment Setup
We conducted experiments on two well-know large-scale text-to-SQL datasets, BIRD (Li et al., 2023b) and Spider (Yu et al., 2018), incorporating multiple robust baselines based on LLMs with in-context learning and fine-tuning. For evaluating text-to-SQL, we utilized Execution Accuracy (EX) (Yu et al., 2018; Li et al., 2023b) and Valid Efficiency Score (VES) (Li et al., 2023b). We included multiple robust baselines from two categories:(i) LLMs + in-context learning baselines: C3-SQL (Dong et al., 2023), DIN-SQL (Pourreza and Rafiei, 2023), MAC-SQL (Wang et al., 2024), DAIL-SQL (Gao et al., 2024), SuperSQL (Li et al., 2024a), Dubo-SQL (Thorpe et al., 2024), TA-SQL (Qu et al., 2024), E-SQL (Caferoğlu and Ulusoy, 2024), CHESS (Talaei et al., 2024); and (ii) LLMs + fine-tuning baselines: DTS-SQL (Pourreza and Rafiei, 2024), CodeS (Li et al., 2024b).
5.2 Main Results of SQL Generation (RQ1)
To address RQ1, we conducted experiments to evaluate the SQL generation capabilities of state-of-the-art (SOTA) baselines enhanced by our KaSLA, with results presented in Table 1. KaSLA replaces the original schema linking component in models that already have schema linking capabilities or implements a full schema prompt substitution for those that do not.
The results demonstrate that incorporating KaSLA consistently improves both Execution Accuracy (EX) and Valid Efficiency Score (VES) across baselines with diverse base models, including GPT-4, CodeS-15B, StarCoder2-15B, and Deepseek-v3. Notably, in the BIRD-dev dataset, KaSLA enhances the EX of CHESS with Deepseek-v3 from 61.34% to 63.30%. For local LLMs, both "CodeS + KaSLA" and "TA-SL + KaSLA" with CodeS-15B and StarCoder-15B achieve impressive increases in Execution Accuracy compared to their original versions. These results underscore KaSLA’s effectiveness and flexibility as a plug-in, showing that KaSLA plays an important role in improving complex text-to-SQL tasks. This adaptability not only confirms its robustness but also broadens its applicability across different models, achieving superior performance outcomes and making it a versatile addition to existing systems.
Base Model Method BIRD-dev Spider-dev
EX VES EX VES
Easy Medium Hard Total Total Easy Medium Hard Extra Total Total
GPT-4 MAC-SQL - - - 57.56 58.76 - - - - 86.75 -
DIN-SQL - - - 50.72 58.79 92.34 87.44 76.44 62.65 82.79 81.70
DAIL-SQL 62.49 43.44 38.19 54.43 55.74 91.53 89.24 77.01 60.24 83.08 83.11
DAIL-SQL (SC) 63.03 45.81 43.06 55.93 57.20 91.53 90.13 75.29 62.65 83.56 -
TA-SQL 63.14 48.82 36.81 56.32 - 93.50 90.80 77.60 64.50 85.00 -
SuperSQL 66.92 46.67 43.75 58.60 60.62 94.35 91.26 83.33 68.67 87.04 85.92
Dubo-SQL - - - 59.71 66.01 - - - - - -
CodeS-15B CodeS 65.19 48.60 38.19 57.63 63.22 94.76 91.26 72.99 64.46 84.72 83.52
CodeS + KaSLA 68.32 50.75 38.19 60.17 64.52 96.37 89.46 76.44 63.86 84.82 84.46
StarCoder2 -15B TA-SL 59.89 45.38 34.72 53.13 60.51 95.56 92.38 78.16 63.25 85.98 85.03
TA-SL + KaSLA 66.59 49.46 38.89 58.80 63.56 94.76 92.15 77.59 66.87 86.27 85.47
DTS-SQL 63.03 46.02 34.72 55.22 64.17 91.94 90.58 78.74 66.87 85.11 85.49
DTS-SQL + KaSLA 63.89 50.11 39.58 57.43 65.20 94.76 90.81 80.46 68.07 86.36 85.81
CodeS 68.00 51.40 39.58 60.30 65.04 94.76 91.26 72.99 64.46 84.72 83.52
CodeS + KaSLA 67.68 54.84 42.36 61.41 66.87 94.76 91.03 78.74 66.27 85.88 84.46
Deepseek -V3 E-SQL 67.24 51.61 41.67 60.10 62.16 95.57 91.93 72.99 64.46 85.21 83.81
E-SQL + KaSLA 69.40 53.55 43.75 62.18 65.32 95.97 92.38 73.56 65.06 85.69 85.72
CHESS 67.89 52.90 46.53 61.34 65.28 95.57 92.16 73.56 65.06 85.50 85.33
CHESS + KaSLA 69.94 54.84 47.92 63.30 67.21 96.37 92.83 74.14 65.66 86.17 86.35
Table 1: The SQL generation performance of enhanced text-to-SQL models using KaSLA, which features 1.6 B parameters combining DeepSeek-coder-1.3B and RoBERTa-Large, is evaluated in terms of Execution Accuracy (EX) (%) and Valid Efficiency Score (VES) (%) on the BIRD-dev and Spider-dev datasets. We utilize KaSLA to substitute the original schema linking component in models that already incorporate schema linking, or to replace the full schema prompt for models that do not.
5.3 Experimental Results of Schema Linking (RQ2)
To answer the RQ2, we conduct the experiments and report the evaluation results of schema linking methods on BIRD-dev in Figure 2. We measured the performance of these schema linking methods in both table linking and column linking. The results indicate that, under the enhanced F1 score which specifically evaluates missing elements, only KaSLA achieves high accuracy in the challenging task of column linking. Schema linking models using LLMs often perform well in table linking but fall short in column linking due to the lack of additional optimizations for missing elements. This highlights the advantages of KaSLA’s framework and demonstrates how enhanced metrics can reveal the significant issue of element missing in current schema linking practices.
Schema linking method Table linking Column linking
DTS-SL (starcoder2-15b) 61.94 23.71
CodeS-SL (starcoder2-15b) 58.21 29.56
RSL-SQL (deepseek-v3) 67.11 17.20
KaSLA (deepseek-coder-1.3B) 81.92 62.78
Table 2: Schema linking performance with the proposed enhanced linking metric on BIRD-dev.
Ablation EX
CodeS + KaSLA 60.17
w/o Binary scoring model 56.06
w/o Probabilistic scoring model 57.82
w/o Hierarchical strategy 58.34
w/o upper limit 1 in Eq. 10 53.26
Table 3: Ablation studies of the key components of KaSLA. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
Dataset Schema linking model Binary scoring function with deepseek-coder-1.3B Probabilistic scoring model with RoBERTa-Large Estimation and dynamic programming Text-to-SQL Total time EX
BIRD-dev Full Schema / / / 5.15 s 5.15 s 57.63
KaSLA 3.5 s 0.12 s $<$ 0.01 s 1.85 s 5.48 s 60.17
Table 4: Inference time cost per instance of each component in KaSLA using CodeS-15B as the text-to-SQL model.
Model SuperSQL CodeS CodeS + KaSLA CodeS + KaSLA (Transfer)
BIRD-dev 58.60 60.30 60.17 60.02
Spider-dev 87.40 84.72 84.82 84.56
Table 5: Execution Accuracy (EX) (%) of KaSLA trained on Spider-train but evaluated on BIRD-dev, and vice versa, for cross-scenario transfer ability evaluation.
5.4 Empirical studies (RQ3)
Ablation Study.
We presented the ablation study results for the key components in KaSLA in Table 3, the results concerning the choice of language model for the probabilistic scoring model are shown in Table 8 in Apendix and the scaling up performance of the scoring model are shown in Table 7 in Apendix, respectively.
As shown in Table 3: (i) Removing the binary scoring model results in a lack of confirmation for high-confidence elements, leading to the inclusion of more redundant ones. This underscores the binary model’s role in filtering out less relevant elements by ensuring that elements with high certainty are correctly identified. (ii) Omitting the probabilistic scoring model risks missing potentially relevant elements, resulting in gaps within schema linking. The probabilistic model is essential for assigning a non-zero probability to all elements, thus capturing those with potential relevance. (iii) The hierarchical strategy effectively reduces the number of candidate columns, thereby lowering complexity and improving efficiency. By structuring the linking process in stages, KaSLA can more precisely target relevant schema components. (iv) Without the upper limit set at 1, the model introduces disproportionate comparisons among similarly relevant items, potentially excluding some relevant elements. This constraint is crucial to maintain balanced relevance assessments, preventing the overshadowing of items that might otherwise contribute positively to SQL generation. These ablation study highlights the significance of key components within KaSLA.
Transferability.
We conducted two experiments to evaluate KaSLA’s performance in cross-scenario transfer. We trained the binary and probabilistic scoring models on the Spider training dataset and evaluated it on the BIRD dev dataset, and vice versa, to explore its cross-scenario transfer ability. Results are provided in Table 5. We can find that pre-training KaSLA on public datasets yields results that outperform the baselines and are only slightly lower than what domain-specific fine-tuning would achieve. The results show that KaSLA demonstrates strong cross-scenario transferability, highlighting its robustness and adaptability across different data domains.
Inference time cost.
We evaluated the inference time cost of integrating KaSLA into text-to-SQL frameworks in Table 4. By utilizing the lightweight LLM DeepSeek-coder-1.3B and significantly reducing the number of tokens in the input prompts through optimized schema linking, KaSLA achieves improved SQL generation accuracy with only a modest increase in inference time. This showcases KaSLA’s efficiency in balancing performance and computational demands.
6 Related Work
Schema linking is essential for aligning natural language queries with database components, playing a pivotal role in the overall text-to-SQL process. Recent approaches utilize large language models (LLMs) to enhance schema linkage through methods such as semantic retrieval, historical query integration, and bidirectional linking (Pourreza and Rafiei, 2024; Cao et al., 2024; Zhang et al., 2024). Techniques like DAIL-SQL and E-SQL refine linkage by leveraging LLMs’ contextual understanding and query enrichment capabilities (Gao et al., 2024; Caferoğlu and Ulusoy, 2024). Retrieval-based strategies focus on retrieve information from database to improve SQL generation (Pourreza and Rafiei, 2024). Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) utilize semantic retrieval to improve schema linking by accessing relevant database information. Despite these advancements, many state-of-the-art text-to-SQL models emphasize fine-tuning strategy. DTS-SQL (Pourreza and Rafiei, 2024) fine-tunes a local LLM to select relevant tables from user queries, while Solid-SQL (Liu et al., 2025) introduces a data augmentation method for enhancing schema linking. DELLM (Hong et al., 2024a) finetuned a specialized data expert LLM to provide additional knowledge for schema linking and SQL generation. Approaches like TA-SQL (Qu et al., 2024) use in-context learning to guide LLMs, generating dummy SQL for linked element abstraction. Similarly, RSL-SQL (Cao et al., 2024) employs a bidirectional approach for schema simplification. Meanwhile, E-SQL (Qu et al., 2024) and (Maamari et al., 2024) note that schema linking challenges persist, leading some models to use the entire schema directly as input for SQL generation.
Different with the previous research, our KaSLA utilize a knapsack optimization framework to avoid the relevant elements missing and the inclusion of redundant elements, optimizing schema linking to enhance SQL generation accuracy effectively.
7 Conclusion
This paper introduced the Knapsack Schema Linking Agent (KaSLA), a pioneering approach to overcoming schema linking challenges in text-to-SQL tasks. Based on knapsack optimization, KaSLA effectively prevent the missing of relevant elements and excess redundancy, thereby significantly enhancing SQL generation accuracy. Our introduction of an enhanced schema linking metric, sets a new standard for evaluating schema linking performance. KaSLA utilizes a hierarchical linking strategy, starting with optimal table linking and then proceeding to column linking within selected tables, effectively reducing the candidate search space. In each step, KaSLA employs a knapsack optimization strategy to link potentially relevant elements while considering a limited tolerance for potential redundancy. Extensive experiments on the Spider and BIRD benchmarks have demonstrated that KaSLA can significantly enhance the SQL generation performance of existing text-to-SQL models by substituting their current schema linking processes. These findings underscore KaSLA’s potential to revolutionize schema linking and advance the broader capabilities of text-to-SQL systems, offering a robust solution that paves the way for more precise and efficient database interactions.
8 Limitation
There are mainly two limitations of this work. First, although KaSLA demonstrates significant advancements in schema linking accuracy and efficiency, its performance traditionally depends on a comprehensive training dataset with detailed ground truth linking results. However, our experiments indicate that KaSLA possesses a degree of transferability, showing promise even in scenarios where training data is sparse or less diverse. This adaptability suggests that KaSLA can maintain reasonable effectiveness across different database environments, although performance might still be impacted in highly dynamic or rapidly evolving schema structures. Additionally, the reliance on a fine-tuning process of LLMs, integrating DeepSeek-coder-1.3B, could pose scalability challenges in resource-constrained settings. Despite these limitations, KaSLA remains a powerful tool for schema linking optimization in text-to-SQL tasks, and future work could focus on enhancing its adaptability and efficiency.
9 Ethics Statement
We confirm that we have fully complied with the ACL Ethics Policy in this study. All the datasets are publicly available and have been extensively used in research related to text-to-SQL.
Appendix A Implementation Details.
We implemented the proposed KaSLA and baselines under the following settings: For top-$K$ used in the redundancy tolerance estimation of KaSLA, we set it to 30. For method using fine-tuned LLM, We utilized LoRA (Hu et al., 2021) for parameter-efficient fine-tuning. The learning rate was initialized to $41e-$ 4 with a cosine decay, the batch size was set to 16, and the training epoch was 3.
Appendix B Input formats
We provided the input and output formats of the binary scoring function of KaSLA introduced in section 4.1 in Table 6.
Type Input format
Full schema table department, columns = [ department__\__id (int |||| primary key||||values: 1), name (text||||values: State, Treasury), ranking (int |||| values: 1, 2)] table head, columns = [ head__\__id (int |||| primary key |||| values: 1), name (text |||| values: Tiger Woods), age (real |||| values: 67.0)] table management, columns = [ department__\__id ( int |||| primary key |||| values: 2), head__\__id (int |||| primary key |||| values: 5), temporary__\__acting (text |||| values: Yes)] foreign keys: management.head__\__id = head.head__\__id, management.department__\__id = department.department__\__id
Question How many heads of the departments are older than 56? Generate SQL to solve the above question and list the relevant columns in each table
Type Output format
Simulated SQL SELECT count(*) FROM head WHERE age > 56
Schema linking results "department": null, "management": null, "head": [ "age" ]
Table 6: Input and output formats of the binary scoring model.
Appendix C The Dynamic Programming Approach used in KaSLA
We illustrate the details of the dynamic programming algorithm used in KaSLA in Algorithm 1.
Input: relevance score set ${\mathcal{R}}$, redundancy score set ${\mathcal{W}}$, the upper redundancy tolerance ${\mathcal{U}}$
Output: selection set $\hat{\mathcal{H}}$
$n\leftarrow|{\mathcal{R}}|$, $^{}^\hat{\mathcal{H}}\leftarrow\{\}$
[…]
$keep\leftarrow\text{array of }(n+1)\times({\mathcal{U}}+1)$
initialized to False
for $i\leftarrow 1$ to $n$ do
       for $w\leftarrow 0$ to ${\mathcal{U}}$ do
             if ${\mathcal{W}}[i-1]\leq w$ then
                   if $({\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]])>\mathcal{A}[i-1][w]$ then
                         $\mathcal{A}[i][w]\leftarrow{\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]]$
                         $keep[i][w]\leftarrow\text{True}$
                  else
                         $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
            else
                   $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
$k\leftarrow{\mathcal{U}}$  for $i\leftarrow n$ downto $1$ do
       if $keep[i][k]$ then
             ^.$\hat{\mathcal{H}}.\text{add}(i-1)$
             $k\leftarrow k-{\mathcal{W}}[i-1]$
return $\hat{\mathcal{H}}$
Algorithm 1 Dynamic Programming-based 0-1 Knapsack Problem Optimization
Appendix D Complexity Analysis
The efficiency of KaSLA stems from its hierarchical approach, dynamic programming and inherent parallelism. KaSLA’s time complexity is determined by two main phases: table linking and column linking within the linked tables. For a database with $n_{t}$ tables and a table capacity of $C_{t}$, the table linking phase has a time complexity of $O(n_{t}C_{t})$. KaSLA then links columns within the selected table. Let $n_{c}$ and $C_{c}$ represent the number of columns and the column capacity of a selected table, respectively. The time complexity of the column linking phase is $O(n_{c}C_{c})$. Consequently, the total time complexity of KaSLA is $O(n_{t}C_{t}+n_{t}(n_{c}C_{c}))$, indicating that the complexity of KaSLA scales relevant with the database dimensions.
Appendix E The Ablation Studies for Different Language Models
We present the ablation studies on the choice of language models for the probabilistic scoring model, as well as the performance scaling of various language models for the binary scoring model used in KaSLA. These studies are detailed in Table 8 and Table 7, respectively.
Dataset BERT -Base BERT -Large RoBERTa -Base RoBERTa -Large SGPT -1.3B
BIRD-dev 59.12 59.52 59.65 60.17 59.97
Table 7: Execution Accuracy (EX) (%) of the scaling up performance about the choice of language model for binary scoring model. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
Dataset Deepseek-coder -1.3B Deepseek-coder -6.5B Deepseek-coder -7B
BIRD-dev 60.17 61.28 61.47
Table 8: Execution Accuracy (EX) (%) of the ablation study about the choice of language model for probabilistic scoring model. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.

### v2

@@
 A crucial reason for this is that commonly used metrics, recall and precision, fail to capture relevant element missing and thus cannot reflect actual schema linking performance.
-Motivated by this, we propose an enhanced schema linking metric by introducing a restricted missing indicator.
-Accordingly, we introduce Knapsack optimization-based Schema Linking Agent (KaSLA), a plug-in schema linking agent designed to prevent the missing of relevant schema elements while minimizing the inclusion of redundant ones.
+Motivated by this, we propose enhanced schema linking metrics by introducing a restricted missing indicator.
+Accordingly, we introduce Knapsack optimization-based Schema Linking Approach (KaSLA), a plug-in schema linking method designed to prevent the missing of relevant schema elements while minimizing the inclusion of redundant ones.
 KaSLA employs a hierarchical linking strategy that first identifies the optimal table linking and subsequently links columns within the selected table to reduce linking candidate space.
-In each linking process, it utilize a knapsack optimization approach to link potentially relevant elements while accounting for a limited tolerance of potential redundant ones.
-With this optimization, KaSLA-1.6B achieves superior schema linking results compared to large-scale LLMs, including deepseek-v3 with state-of-the-art (SOTA) schema linking method.
-Extensive experiments on Spider and BIRD benchmarks verify that KaSLA can significantly improve the SQL generation performance of SOTA text-to-SQL models by substituting their schema linking processes.
-Knapsack Optimization-based Schema Linking for
-LLM-based Text-to-SQL Generation
-Zheng Yuan1, Hao Chen2, Zijin Hong1, Qinggang Zhang1 Feiran Huang3, Xiao Huang1 1The Hong Kong Polytechnic University 2City University of Macau, 3Jinan University {yzheng.yuan, zijin.hong, qinggangg.zhang}@connect.polyu.hk sundaychenhao@gmail.com; huangfr@jnu.edu.cn; xiao.huang@polyu.edu.hk
-1 Introduction
-Figure 1: In text-to-SQL tasks, commonly used schema linking metrics such as Recall and Precision fail to accurately capture the true performance of schema linking when relevant schema elements are missing.
-Figure 2: Overall framework of proposed knapsack optimization-based schema linking agent (KaSLA) with column linking as an example.
-Traditional schema linking models often struggle with both missing relevant elements and the inclusion of redundant ones.
-KaSLA addresses this challenge by maximizing the total estimated relevance while maintaining total redundancy within a specified tolerance using knapsack optimization.
-Serving as a general model, KaSLA enhances the SQL generation of text-to-SQL models by replacing their existing schema linking processes.
-With the advent of large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024), LLM-based text-to-SQL is emerging as the next-generation interface for database users (Hong et al., 2024b).
-Typically, text-to-SQL frameworks employ a two-step process: first linking database schema elements (tables and columns) to user queries, then generating the corresponding SQL statements.
-However, the accurate SQL generation requires the linking to the correct schema elements, meaning that errors in early stages are inevitably propagated to later SQL generation.
-For instance, inaccurate schema linking—such as overlooking relevant schema elements or linking to irrelevant ones—will lead to incorrect SQL statements.
-Recent state-of-the-art text-to-SQL models tend to focus primarily on final SQL generation while employing relatively simplistic schema linking strategies or neglecting it.
-DTS-SQL (Pourreza and Rafiei, 2024) finetunes a local LLM to select relevant tables based on user queries.
-Solid-SQL (Liu et al., 2025) present a data augmentation method for the finetuning of Schema linking method.
-Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) use semantic-based retrieval models to enhance the schema linking process by retrieving relevant information from databases.
-RSL-SQL (Cao et al., 2024) introduces a bidirectional approach to simplify the full schema.
-E-SQL (Qu et al., 2024) and (Maamari et al., 2024) assert that schema linking remains challenging for current LLMs and thus directly use the full schema as input for SQL generation.
-Even though schema linking is crucial, current widely used metrics, such as Precision and Recall (Li et al., 2023a, 2024b; Qu et al., 2024), cannot adequately evaluate instances when relevant elements are missing.
-As demonstrated in Figure 1, missing a single relevant table or column can significantly decrease the accuracy of SQL generation, yet still result in high Recall and Precision scores.
-Because Recall does not significantly penalize the model as long as the majority of the relevant elements are identified, allowing it to achieve an artificially high score, even when crucial elements are missing.
-Similarly, regardless of the severity of the missing, as long as the number of correctly linked elements is proportionally high compared to the total predicted relevant elements, the model can achieve an unreasonably high Precision score.
-This discrepancy distorts the assessment of a schema linking model’s effectiveness, potentially hindering progress in text-to-SQL development.
-To address this, we firstly design a tailored schema linking metrics that accurately assess linking performance with a restricted element missing indicator.
-Additionally, we propose the Knapsack optimization-based Schema Linking Agent (KaSLA), designed to link the most relevant and minimally redundant schema elements according to the given queries.
-KaSLA introduces a binary-probabilistic score function to predict a robust relevance score of each element.
-Additionally, KaSLA identifies redundancy of each element and the upper redundancy tolerance for each query based on its most similar queries in the training dataset, effectively minimizing the inclusion of redundant elements.
-Finally, KaSLA links tables and columns with an hierarchical strategy to reduce the column linking candidate space for efficiency and employs a 0-1 knapsack optimization approach to select the most relevant element set under the upper redundancy tolerance to reduce the redundant linking.
-Extensive experiments demonstrate that KaSLA can improve SQL generation accuracy by reducing missing and redundant information in the schema linking.
+In each linking process, it utilizes a knapsack optimization approach to link potentially relevant elements while accounting for a limited tolerance of potentially redundant ones.
+With this optimization, KaSLA-1.6B achieves superior schema linking results compared to large-scale LLMs, including deepseek-v3 with the state-of-the-art (SOTA) schema linking method.
+Extensive experiments on Spider and BIRD benchmarks verify that KaSLA can significantly improve the SQL generation performance of SOTA Text2SQL models by substituting their schema linking processes.
+Index Terms:
+text-to-SQL, database, large language models, natural language understanding
+I Introduction
+With the advent of large language models (LLMs) [1, 2], LLM-based text-to-SQL is emerging as the next-generation database interface for unskilled users [3].
+Typically, text-to-SQL frameworks employ a two-stage process: first linking user queries to database schema elements (tables and columns), then generating the corresponding SQL statements.
+Accurate SQL generation relies on correctly identifying these schema elements, since errors in this initial linking phase inevitably propagate to subsequent SQL construction.
+Inaccurate schema linking, such as overlooking relevant schema elements or linking to irrelevant ones, consistently produces erroneous SQL.
+This challenge is shown in Figure 1.
+When using CodeS-15B, a pre-trained LLM for SQL generation, on the BIRD-dev benchmark, there is a significant 14.91% performance gap between state-of-the-art schema linking methods and ground-truth linking results.
+Consequently, enhancing schema linking accuracy represents a critical research frontier with significant potential to advance text-to-SQL capabilities.
+Figure 1: Performance comparison of different schema linking methods on the BIRD-dev dataset, using CodeS-15B as the SQL generation model.
+Figure 2: Commonly used metrics, like Recall and Precision, fail to accurately reflect actual schema linking performance when relevant elements are missing.
+Recent state-of-the-art text-to-SQL models primarily focus on the final SQL generation stage, often relying on basic or even neglecting schema linking strategies.
+Existing schema linking methods can generally be categorized into three types.
+First, some frameworks directly use the full database schema as input without a specialized schema linking mechanism, instead depending on the reasoning capabilities of LLMs.
+For example, DAIL-SQL [4], E-SQL [5], and Distillery-SQL [6] enhance LLMs’ ability to process full schema through effective prompting.
+Second, recall-based schema linking methods have been developed.
+For instance, ResdSQL [7] and CodeS [8] train encoder-based language models to rank and predict relevant schema elements.
+They employ decoupled linking strategies to separately handle tables and columns, enabling more effective use of semantic information.
+Third, generative schema linking methods typically leverage LLMs as schema linking models.
+DTS-SQL [9] finetunes a local LLM to select relevant tables given user queries, while Solid-SQL [10] proposes data augmentation to improve schema linking performance.
+Other approaches, such as TA-SQL [5], use in-context learning to guide LLMs in generating dummy SQL that abstracts linked schema elements.
+Multi-agent models like CHESS [11] design dedicated schema selector agents that prompt LLMs to choose relevant tables and columns.
+Furthermore, RSL-SQL [12] adopts a bidirectional approach for schema simplification, aiming to further refine schema linking.
+Despite the critical role of schema linking, traditional metrics like Recall and Precision [7, 8, 5] fail to assess missing relevant elements adequately.
+A single missing element can render SQL generation impossible, yet current metrics may still yield high scores that misrepresent true model performance.
+As shown in Figure 2, the schema linking result $\widehat{{\bm{S}}}_{q}$ misses one key column, organization, which directly causes the failure to generate correct SQL due to incomplete schema linking results.
+However, traditional Recall and Precision metrics still give $\widehat{{\bm{S}}}_{q}$ falsely high scores.
+This happens because Recall does not severely penalize the model if most relevant elements are found, allowing a high score despite missing a few relevant elements.
+Likewise, Precision can be unreasonably high as long as the number of correctly linked elements is large compared to the total predicted ones, regardless of how many are missed.
+This discrepancy skews schema linking evaluation and may hinder progress in text-to-SQL research.
+To address this issue, we propose tailored schema linking metrics, Recall+ and Precision+, which incorporate a restricted missing-element indicator to more accurately assess linking performance.
+These metrics impose penalties when relevant elements are missing, since the absence of any relevant element in the schema linking results prevents the SQL generation model from obtaining the complete information needed to generate the correct SQL.
+Based on the proposed evaluation metrics, we find that existing schema linking methods face a significant issue—the Missing & Redundancy Seesaw Problem: There exists a seesaw phenomenon between missing and redundant elements in current schema linking methods: Generative linking strategies tend to have fewer redundant elements but more missing ones, while recall-based models usually have fewer missing elements but more redundancy.
+It is difficult for current methods to achieve both low missing and low redundancy at the same time, creating a trade-off situation similar to a seesaw.
+Developing schema linking approaches that can effectively reduce both missing and redundant schema elements remains a significant challenge.
+To address this seesaw problem, we propose the Knapsack Optimization-based Schema Linking Aapproach (KaSLA), aims to link the relevant schema elements for each query while minimizing missing and redundant cases simultaneously.
+KaSLA introduces a binary-probabilistic score function to predict a robust relevance score for each element.
+In addition, KaSLA identifies redundancy of each element and determines an upper redundancy tolerance for each query by referencing its most similar queries in a pre-constructed pool of query and linking result pairs.
+Finally, KaSLA performs table and column linking hierarchically; it links columns based on intermediate table linking results, which reduces the candidate search space.
+During each linking step, KaSLA uses a 0-1 knapsack optimization approach to select the most relevant set of elements under the defined upper redundancy tolerance, aiming to avoid missing elements and minimize redundancy, all within an efficient dynamic programming framework.
+Extensive experiments show that KaSLA improves a text-to-SQL framework’s SQL generation accuracy by simultaneously reducing missing and redundant information in schema linking.
 In summary, our contributions are as follows:
-  • We introduce an enhanced schema linking metric, which can accurately capture the missing of relevant elements and reflect the actual linking performance aligning with the SQL generation performance.
-  • We propose the Knapsack optimization-based Schema Linking Agent (KaSLA) to ensure the linking of relevant schema elements and minimizing the inclusion of redundant ones.
-  • KaSLA can serve as a general agent to enhance LLM-based text-to-SQL framework by provide optimal linking results.
+  • We introduce enhanced schema linking metrics, Recall+ and Precision+, which can accurately capture the missing of relevant elements and reflect the actual linking performance aligning with the SQL generation.
+  • We propose the Knapsack optimization-based Schema Linking Aapproach (KaSLA) to link the relevant schema elements for each query while minimizing missing and redundant elements simultaneously.
+  • KaSLA can serve as a general plug-in method to enhance LLM-based text-to-SQL framework by provide optimal linking results.
 Extensive experiments demonstrate improved performance in the SOTA text-to-SQL models when using KaSLA for schema linking.
-2 Preliminaries
-In this section, we outline the definitions for the schema linking and 0-1 Knapsack Optimization.
-Schema Linking Task.
+II Preliminary Study
+In this section, we outline the definitions for the schema linking task, SQL generation task and 0-1 knapsack optimization.
+Additionally, we introduce the traditional schema linking metrics and then discuss their limitations.
+II-A Definitions
+II-A1 Schema Linking Task
 Given a natural language query $q$ and a corresponding database schema ${\mathcal{S}}_{q}$ containing schema elements, tables ${\mathcal{T}}=\{t_{1},\cdots,t_{|{\mathcal{T}}|}\}$ and columns ${\mathcal{C}}=\{c_{1},\cdots,c_{|{\mathcal{C}}|}\}$ in each table.
-Schema linking is a common process used in text-to-SQL to predict the relevant schema entities with $q$ from ${\mathcal{S}}_{q}$:
+Schema linking is a common process used in text-to-SQL to predict the relevant schema elements (tables and columns) with $q$ from ${\mathcal{S}}_{q}$:
 $f_{\text{linking}}:(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q},$ (1)
-where $\widehat{{\bm{S}}}_{q}$ is the subset of ${\mathcal{S}}_{q}$ that includes schema element $s\in{\mathcal{S}}_{q}$ that relevant to $q$.
+where $\widehat{{\bm{S}}}_{q}$ is a subset of ${\mathcal{S}}_{q}$ that includes schema element $s\in{\mathcal{S}}_{q}$ that relevant to $q$ and will be used for subsequent SQL generation.
 Each $s$ could represent a specific table or column.
 Additionally, we denote ${\bm{S}}_{q}^{*}$ as the ground truth linking results.
-0-1 Knapsack Optimization.
-The 0-1 knapsack problem Fréville (2004) is a classic combinatorial optimization challenge.
+II-A2 SQL Generation Task
+With the natural language query $q$ and the schema linking results $\widehat{{\bm{S}}}_{q}$, the function of SQL generation task is to generate an SQL statement that is able to retrieve the pertinent information of $q$ from the database:
+$f_{\text{sql}}:(q,\widehat{{\bm{S}}}_{q})\mapsto{\text{S}\widehat{\text{Q}}\text{L}}_{q},$ (2)
+where $\text{S}\widehat{\text{Q}}\text{L}_{q}$ represents the predicted SQL statement.
+II-A3 0-1 Knapsack Optimization
+The 0-1 knapsack problem [13] is a classic combinatorial optimization challenge.
 It involves selecting items from a finite item set, each with a value ${\mathcal{V}}$ and weight ${\mathcal{W}}$, to place into a knapsack with a maximum weight capacity ${\mathcal{C}}$.
@@
 where $\widehat{N}$ is the selected item set for the knapsack.
-3 Enhanced Schema Linking Metrics
-In this section, we introduce enhanced schema linking metrics considering the missing of relevant elements during evaluation, thereby accurately assess the true performance of schema linking models.
-For a query $q$ with predicted schema linking result $\widehat{{\bm{S}}}_{q}$ and ground truth linking result ${\bm{S}}^{*}_{q}$, Recall and Precision are two commonly used metreics (Li et al., 2023a, 2024b; Qu et al., 2024).
-Recall, ${\bm{R}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|{\bm{S}}^{*}_{q}|}$, measures the proportion of relevant elements correctly identified, specifically how many of the actual relevant elements in ${\bm{S}}^{*}_{q}$ are found in $\widehat{{\bm{S}}}_{q}$.
-Conversely, Precision, ${\bm{P}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|\widehat{{\bm{S}}}_{q}|}$, assesses the accuracy of predictions by determining the proportion of elements predicted as relevant in $\widehat{{\bm{S}}}_{q}$ that are indeed correct.
-As illustrated in Figure 1, these metrics can yield high scores even in the missing of relevant elements.
-This underscores that they do not effectively address the issue of missing relevant elements, which can undermine the true accuracy of schema linking models.
+II-B Traditional Schema Linking Metrics
+II-B1 AUC
+AUC is utilized to evaluate schema linking by treating it as a binary classification task, relevant elements are labeled as 1 and irrelevant ones as 0 [5].
+However, as depicted in Figure 3, the significant imbalance between relevant and irrelevant elements will hinder the AUC from accurately reflecting the linking performance.
+A model can cheat its way to a high AUC by classifying many irrelevant elements, an easier task than accurately identifying the few relevant ones.
+II-B2 Recall and Precision
+Recall and Precision are commonly used in recent text-to-SQL research to evaluate schema linking performance [7, 8, 5].
+For a query $q$ with the predicted schema linking result $\widehat{{\bm{S}}}_{q}$, their formula is shown below:
+${\bm{R}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|{\bm{S}}^{*}_{q}|},{\bm{P}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|\widehat{{\bm{S}}}_{q}|},$ (4)
+where ${\bm{S}}^{*}_{q}$ is the ground truth linking result.
+Recall ${\bm{R}}(q)$ measures the proportion of relevant elements correctly identified, specifically how many of the actual relevant elements in ${\bm{S}}^{*}_{q}$ are found in $\widehat{{\bm{S}}}_{q}$.
+Conversely, Precision ${\bm{P}}(q)$ assesses the accuracy of predictions by determining the proportion of elements predicted as relevant in $\widehat{{\bm{S}}}_{q}$ that are indeed correct.
+As shown in Figure 2, the predicted schema linking result $\widehat{{\bm{S}}}_{q}$ fails to include an essential column, “organization”.
+Without this information, the SQL generation model cannot generate the correct SQL statement.
+Nonetheless, Recall and Precision still report high scores, as they mainly reflect the number of correctly identified elements but overlook the impact of missing relevant ones.
+This demonstrates that these metrics can overestimate the effectiveness of schema linking, especially in this scenario where even a single omission is detrimental.
+Figure 3: The imbalanced proportion of relevant schema elements with the irrelevant ones in the commonly used text-to-SQL datasets.
+Figure 4: Missing & redundancy seesaw problem in schema linking. We evaluate the column linking performance of schema linking baselines on BIRD-dev.
+III Enhanced Schema Linking Metrics
+In this section, we introduce enhanced schema linking metrics considering the absence of relevant elements during evaluation, thereby accurately assessing the performance of schema linking models.
+Addtionaly, we discuss the missing & redundancy seesaw problem in current schema linking methods.
+III-A Recall+ and Precision+:
 To address the limitations of Recall and Precision, we firstly introduce a restricted element missing indicator to detect whether relevant elements are missing in the predicted schema linking result $\widehat{{\bm{S}}}_{q}$, The formal definition is as follows:
@@
 Conversely, it is set to 0 when there are any missing.
-Building on this missing indicator, we define enhanced Recall and Precision as follows:
+Building on this missing indicator, we define Recall+ and Precision+ as follows:
 ${\bm{R}}^{\text{{+}}}(q)=\bm{\mathbbm{1}}(q){\bm{R}}(q),{\bm{P}}^{\text{{+}}}(q)=\bm{\mathbbm{1}}(q){\bm{P}}(q).$ (6)
-Such metrics strengthen the penalty for missing relevant elements, as schema linking results with missing are inherently incorrect for SQL generation.
-For simplicity, we use the harmonic mean of enhanced Recall and Precision, referred to as the enhanced F1 score, as the overall measure for evaluating schema linking results:
+Such metrics strengthen the penalty for missing any relevant element, as schema linking results with missing are inherently incorrect for SQL generation.
+Figure 5: Overall framework of proposed knapsack optimization-based schema linking approach (KaSLA) and and how it operates within a text-to-SQL system.
+Traditional schema linking models often struggle with both missing relevant elements and the inclusion of redundant ones.
+KaSLA addresses this challenge by maximizing the total estimated relevance while maintaining total redundancy within a specified tolerance using knapsack optimization.
+Serving as a general plug-in method, KaSLA enhances the SQL generation of text-to-SQL models by replacing their existing schema linking processes.
+Since Recall+ evaluates whether there are missing elements in the linking results and Precision+ measures the proportion of redundant elements, we use their harmonic mean, called F1+, as the overall metric.
+The formula is as follows:
 ${{\bm{F}}1}^{\text{{+}}}(q)=\frac{2{\bm{R}}^{\text{{+}}}(q){\bm{P}}^{\text{{+}}}(q)}{{\bm{R}}^{\text{{+}}}(q)+{\bm{P}}^{\text{{+}}}(q)}.$ (7)
-This approach provides a comprehensive evaluation of both recall and precision while incorporating the critical aspect of missing elements.
-4 Knapsack Optimization-based Schema Linking Agent (KaSLA)
-In this section, we introduce the proposed KaSLA framework, which focuses on ensuring the linking of relevant schema elements to avoid missing and minimizing the inclusion of redundant ones.
-As illustrated in Figure 2, We begin by detailing a binary-probabilistic score function used for assessing the relevance of schema elements and an estimation function for the redundancy tolerance of each schema linking process.
-We then outline the knapsack optimization approach to link potential relevant element while reducing the potential redundant ones under the tolerance constraint.
-Finally, we describe the hierarchical schema linking process, demonstrating its efficiency for reducing the column linking candidate.
-4.1 Relevance Estimation
+F1+ combines the strengths of Recall+ and Precision+ and provides a comprehensive evaluation of schema linking.
+Taking the same example in Figure 2, our enhanced metrics Recall+ and Precision+ offer a more reliable evaluation.
+When the predicted result lacks any key element, such as the “organization” column, these metrics assign a score of zero, signaling a definite failure in schema linking for SQL generation.
+By strictly penalizing missing crucial information, Recall+ and Precision+ ensure models are only rewarded for comprehensive and complete predictions, thus aligning the metric outcomes closely with actual SQL generation success.
+III-B The Missing & Redundancy Seesaw Problem
+After eliminating the misleading effects caused by missing elements, Recall+ and Precision+ can be used to accurately measure a schema linking model’s ability to link relevant elements and avoid redundant elements, respectively.
+We applied several baselines on BIRD-dev and evaluated their column linking preferences.
+As shown in Figure 4, we observed a significant seesaw phenomenon between missing and redundancy in these schema linking methods.
+For example, recall-based models like CodeS-SL rarely miss relevant columns, resulting in high Recall+, but it includes too many redundant columns, leading to very low Precision+.
+In contrast, generative methods such as Pure-SL, RSL-SL and TA-SL tend to have more missing columns, resulting in lower Recall+, but their linking results contain fewer redundant columns, yielding relatively higher Precision+ compared to the recall-based model, CodeS-SL.
+This prominent missing & redundancy seesaw problem inspires a key question in schema linking: How can we minimize both element missing and element redundancy while ensuring that relevant elements are properly linked?
+To address this issue, we propose the Knapsack Optimization-based Schema Linking Approach (KaSLA), which uses a knapsack optimization framework to improve the schema linking process.
+The details are provided in Section IV.
+IV Knapsack Optimization-based Schema Linking Approach (KaSLA)
+In this section, we present KaSLA, a novel schema linking framework for text-to-SQL tasks that explicitly addresses both missing and redundant element challenges through a knapsack optimization approach.
+We first provide an overview of the KaSLA framework in Section IV-A, outlining its key components, optimization objective and hierarchical structure.
+Next, we introduce the core components in Section IV-B, including the hybrid relevance estimation mechanism, redundancy assessment strategy, and the formulation of redundancy tolerance for each query.
+We then describe how schema linking works as a knapsack optimization process and explain the hierarchical linking strategy for efficient and precise schema element selection in Section IV-C.
+Each subsection details a critical aspect of KaSLA, jointly enabling accurate and efficient schema linking for robust text-to-SQL generation.
+IV-A Overall Framework
+As illustrated in Figure 5, KaSLA is designed to systematically address the challenges of both missing and redundant schema elements in text-to-SQL systems by using a knapsack optimization approach.
+The overall process begins with two key inputs: a natural language query and the full database schema.
+As visible in the left portion of Figure 5, KaSLA follows a hierarchical linking strategy, where it first performs table linking to identify tables relevant to the query, and then proceeds to link columns within each selected table.
+This approach effectively reduces the candidate space and enhances efficiency when working with complex database schema.
+At the core of KaSLA’s schema linking is a hybrid scoring module that combines a binary scoring function (yielding hard 0/1 relevance decisions) and a probabilistic scoring function (providing soft relevance probabilities) for each schema element.
+As shown in the right side of Figure 5, these scores are integrated to assign a final relevance score $r$ to every candidate element.
+In parallel, each element receives a redundancy score $w$, and for each query, a redundancy tolerance $u$ is estimated by analyzing similar queries with their linking results from a pre-constructed query-linking pool, as depicted in the figure.
+This step ensures the model adapts its tolerance to different query types and schema complexities.
+The selection of relevant schema elements is then formulated as a 0-1 knapsack problem: KaSLA seeks to maximize the total relevance scores of linked elements while keeping the sum of their redundancy scores within the estimated tolerance $u$.
+The optimization objective, shown in red in Figure 5, is solved efficiently using a dynamic programming algorithm.
+The knapsack selection process ensures that the final set of linked schema elements contains all necessary information for SQL generation while minimizing unnecessary redundancy.
+After schema linking, the selected elements are passed to the downstream SQL generation model, which constructs the final SQL statement in accordance with the linking result.
+As shown on bottom left of Figure 5, this process allows KaSLA to serve as a plug-in method for existing text-to-SQL frameworks, helping them avoid common pitfalls of missing relevant elements or introducing redundant ones.
+By leveraging hierarchical optimization, hybrid scoring and redundancy-aware selection, KaSLA robustly improves schema linking across diverse and large-scale text-to-SQL applications.
+IV-B Key Factor Estimation
+In this section, we present a detailed introduction to the estimation functions of key factor in KaSLA: element relevance, element redundancy and upper redundancy tolerance.
+IV-B1 Relevance Estimation
 To precisely assess the relevance of each schema element $s$ in relation to the given query $q$, we employ a hybrid binary-probabilistic score function.
 This function integrates a generative model for binary scoring alongside an encoding model for probabilistic scoring, allowing for a comprehensive estimation of relevance.
-Binary scoring function.
-This component is tasked with providing binary relevance assessments, determining whether a schema element is relevant or not in a straightforward manner:
+Binary Scoring Function
+This component is tasked with providing binary relevance assessments, determining whether a schema element is relevant or not in a straightforward manner with hard 0/1 relevance decisions:
 $f_{\text{binary}}:(q,{\mathcal{S}}_{q})\mapsto\{r^{b}_{1},\cdots,r^{b}_{|{\mathcal{S}}_{q}|}\},$ (8)
 where $r^{b}_{i}\in\{0,1\}$ represents the binary score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$. A score of 1 indicates relevant while 0 signifies redundant.
-Considering efficiency, we finetune a lightweight LLM, DeepSeek-Coder-1.3B (Guo et al., 2024), using LoRA (Hu et al., 2021), a well-established parameter-efficient technique.
+Considering efficiency, we finetune a lightweight LLM, DeepSeek-Coder-1.3B [14], with a well-established parameter-efficient technique, LoRA [15].
 We employ a strategy whereby the initial generation of a simulated SQL provides a contextual foundation for the subsequent generation of schema linking.
@@
 Noted that the simulated SQL generated during this process will not be utilized in subsequent steps.
-The input and output examples are provided in Table 6 in Appendix B.
-Probabilistic scoring function.
+Probabilistic Scoring Function
 In conjunction with the binary scoring model, we introduce a probabilistic scoring model to enhance prediction robustness.
 The primary motivation is to mitigate the risk of the binary model misclassifing relevant elements.
-The probabilistic scoring model assigns a soft score to each element as the linking probability to reduce the likelihood of relevant elements missing , thereby supporting the binary model:
+The probabilistic scoring model assigns a soft score to each element as the linking probability to reduce the likelihood of the missing of potential relevant elements, thereby supporting the binary model.
+The equation is defined as follows:
 $f_{\text{prob.}}:(q,{\mathcal{S}}_{q})\mapsto\{r^{p}_{1},\cdots,r^{p}_{|{\mathcal{S}}_{q}|}\},$ (10)
 where $r^{p}_{i}\in[0,1]$ represents the probabilistic score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$.
-For the probabilistic scoring model, we utilize an encoding model, RoBERTa-Large (Liu et al., 2019), to derive the semantic embeddings for the query and all schema elements.
-Following the methodology of (Li et al., 2023a, 2024b), we employ a cross-attention network to jointly embed the semantic representations of tables and their columns.
+A value closer to 1 indicates that the element is considered more relevant to query q.
+For the probabilistic scoring model, we utilize an encoding model, RoBERTa-Large [16], to derive the semantic embeddings for the query and all schema elements.
+Following the methodology of  [7, 8], we employ a cross-attention network to jointly embed the semantic representations of tables and their columns.
 The dot product of the query embedding and the element embedding yields the final probabilistic score.
@@
 $\mathcal{L}_{\text{prob.}}=\sum_{q\in Q}\text{FL}(\mathbbm{1}(s\in{\bm{S}}^{*}_{q}),f_{\text{prob.}}(q,{\mathcal{S}}_{q})),$ (11)
-where $\text{FL}(\cdot)$ denotes the focal loss function (Ross and Dollár, 2017), designed to emphasize learning from hard negative samples.
-Relevance scoring estimation.
+where $\text{FL}(\cdot)$ denotes the focal loss function  [17], designed to emphasize learning from hard negative samples.
+Relevance Scoring Estimation
 The final function for estimating the relevance of each schema element integrates the scores derived from both the binary and probabilistic functions:
@@
 An upper limit of 1 is imposed on $r_{i}$ to prevent excessively strong contrasts between potentially relevant elements, as each relevant schema element is equally important for accurate SQL generation.
-4.2 Redundancy Estimation
+We use ${{\mathcal{R}}}_{{\mathcal{S}}_{q}}=\{r_{1},\ldots,r_{|{\mathcal{S}}_{q}|}\}$ to denote the relevance score set of ${\mathcal{S}}_{q}$.
+IV-B2 Redundancy Estimation
 To minimize the inclusion of redundant elements during schema linking, we define both a redundancy score for each schema element $s\in{\mathcal{S}}_{q}$ and the upper redundancy tolerance for query $q$.
-Redundancy scoring estimation.
+Redundancy Scoring Estimation
 Recognizing that elements with lower relevance scores are more likely to be redundant, whereas those with higher scores are more essential, we assign a redundancy score $w_{i}$ to each schema element $s_{i}\in{\mathcal{S}}_{q}$ in relation to query $q$.
 This score is calculated as the inverse of its relevance score $r_{i}$, formulated as follows:
-$w_{i}={r_{i}}^{-1}.$ (11)
+$\begin{split}f_{\text{redundancy}}:&(q,{{\mathcal{R}}}_{{\mathcal{S}}_{q}})\mapsto\{w_{1},\ldots,w_{|{\mathcal{S}}_{q}|}\}\\ &=\left\{{r_{i}}^{-1}\mid r_{i}\in{{\mathcal{R}}}_{{\mathcal{S}}_{q}}\right\}.\end{split}$ (13)
 It ensures elements with lower relevance scores contribute more to redundancy, thereby effectively prioritizing more relevant schema elements.
-Upper Redundancy Tolerance.
-Redundancy tolerance defines the limit on the total redundancy score of elements in the schema linking results, as an excess of redundant elements can significantly disrupt subsequent SQL generation.
-Estimating it during inference is challenging.
+We use ${{\mathcal{W}}}_{{\mathcal{S}}_{q}}=\{w_{1},\ldots,w_{|{\mathcal{S}}_{q}|}\}$ to denote the redundancy score set of ${\mathcal{S}}_{q}$.
+Redundancy Tolerance
+Upper tolerance of redundancy defines the limit on the total redundancy score of elements in the schema linking results, as an excess of redundant elements can significantly disrupt subsequent SQL generation.
+Estimating a suitable upper redundancy tolerance during inference is challenging.
 For the SQL generation of some queries, a lenient tolerance may be beneficial since it allowing the inclusion of more potentially relevant elements in the linking results to prevent missing.
-However, for other queries, a lenient tolerance could introduce to many redundant elements, which may significantly disrupt the final SQL generation process.
-To address this challenge, we hypothesize that similar queries exhibit comparable tolerance levels for redundancy and leverage queries from the training dataset, where ground truth linking results are available, to aid in predicting redundancy tolerance for given query $q$ during inference.
-We first retrieve the top-$K$ queries in the training dataset that closely resemble $q$ using sentence similarity measures (Gao et al., 2021).
-For a query $q_{k}$ in the retrieved top-$K$ queries, we calculate the total redundancy score for its ground truth linking results ${\mathcal{S}}^{*}_{q_{k}}$ and use it as its upper redundancy tolerance.
-The maximum of these upper redundancy tolerances from the $K$ most similar queries in the training dataset is then used as the redundancy tolerance estimate for query $q$:
-$u_{q}=\max_{q_{k}\in\text{TopK}(q)}\bigg{(}\sum_{s_{j}\in{\mathcal{S}}^{*}_{q_{k}}}w_{j}\bigg{)}$ (12)
-By capturing the maximum sum from the most similar queries, this approach effectively defines an upper bound on redundancy tolerance that is both efficient and adaptable, ensuring that while redundant elements are excluded, the risk of missing relevant elements is minimized.
-4.3 Schema Linking with Knapsack Optimization
-Given the relevance score $r_{i}$, the redundancy score $w_{i}$ for each $s_{i}\in{\mathcal{S}}_{q}$, and the upper redundancy tolerance $u_{q}$ for query $q$, we introduce a hierarchical schema linking strategy with knapsack optimization.
+However, for other queries, a lenient tolerance could introduce too many redundant elements, which might significantly disrupt the final SQL generation process.
+To address this challenge, we hypothesize that similar queries have comparable tolerance levels for redundancy.
+We sample queries from the training dataset where ground truth linking results are available, and use these paired samples to construct a query-linking pool.
+During inference, when predicting the redundancy tolerance for a given query $q$, we first retrieve the top-$K$ most similar queries from the query-linking pool using sentence similarity measures [18].
+For a query $q_{k}$ in the retrieved top-$K$ queries, the redundancy score set for its ground truth linking results ${\mathcal{S}}^{*}_{q_{k}}$ can be denoted as ${{\mathcal{W}}}_{{\mathcal{S}}^{*}_{q_{k}}}$.
+We sum the total ${{\mathcal{W}}}_{{\mathcal{S}}^{*}_{q_{k}}}$ and use it as the upper redundancy tolerance of $q_{k}$.
+The maximum of these upper redundancy tolerances of the $K$ most similar queries from the query-linking pool is then used as the redundancy tolerance estimate for query $q$:
+$f_{\text{tolerance}}:(q)\mapsto u_{q}=\max_{q_{k}\in\text{TopK}(q)}\sum_{w_{j}\in{{\mathcal{W}}}_{{\mathcal{S}}^{*}_{q_{k}}}}w_{j}.$ (14)
+By capturing the maximum redundancy tolerances of the most similar queries, this approach effectively defines an upper bound on redundancy tolerance that is both efficient and adaptable, ensuring that while redundant elements are excluded, the risk of missing relevant elements is minimized.
+IV-C Schema Linking with Knapsack Optimization
+Given (i) the relevance score $r_{i}$ and redundancy score $w_{i}$ for each candidate element $s_{i}\in{\mathcal{S}}_{q}$, and (ii) the upper redundancy tolerance $u_{q}$ defined at the query level, we propose a hierarchical schema linking strategy incorporating knapsack optimization.
 This approach is designed to ensure the linking of relevant schema elements to avoid missing and minimize the inclusion of redundant ones for a precise schema linking.
-Knapsack Optimization
-The 0-1 knapsack problem (Fréville, 2004) is a classic combinatorial optimization challenge.
-It involves selecting items from a finite set, each item characterized by a specific value and weight, to place into a knapsack with a limited weight capacity.
-The objective of 0-1 knapsack optimization is to choose items that maximize the total value while ensuring the total weight remains below the knapsack’s capacity.
+IV-C1 Knapsack Optimization
+The 0-1 knapsack problem [13] is a classic combinatorial optimization problem.
+It involves selecting items from a finite set.
+Given a set of items, each with a value and a weight, the goal is to select a subset of items that maximizes the total value while ensuring the total weight does not exceed a limited knapsack capacity.
+In the “0-1” variant, each item is either fully included or excluded (i.e., no fractional selections are permitted).
 Inspired by this, we develop a knapsack optimization-based schema linking strategy to ensure the inclusion of relevant schema elements while minimizing redundant ones, thus ach ieving optimal schema linking.
 Formally, the optimization objective are defined as follows:
-$\begin{split}f_{\text{Knap}}:&(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q}=\arg\max\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}r_{i}\\ &\text{s.t.}\quad\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}w_{i}\leq u_{q}.\end{split}$ (13)
-This formulation aims to maximize the total relevance scores of the linked schema elements.
+$\begin{split}f_{\text{Knap}}:&(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q}\\ &=\arg\max\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}r_{i},\quad\text{s.t.}\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}w_{i}\leq u_{q}.\end{split}$ (15)
+This formulation aims to maximize the total relevance scores of the linked schema elements with a total redundancy constraint.
 The constraint ensures that the cumulative redundancy score remains within the upper redundancy tolerance of query $q$.
-To tackle this optimization problem and achieve the optimal linking result, We employ an efficient dynamic programming approach.
-Detailed pseudocode of this dynamic programming algorithm is provided in Algorithm A in Appendix.
-Hierarchical Linking Strategy
-For efficiency, our KaSLA employs a hierarchical process in schema linking: initially linking tables and subsequently linking columns within each selected table.
-This strategy effectively reduces the dimensionality of candidate columns, enhancing KaSLA’s capacity to handle large-scale schema linking and text-to-SQL generation tasks.
-In the initial phase for a given query $q$, KaSLA aims to identify the relevant tables from the set of all available tables in schema ${\mathcal{S}}_{q}$, denoted as ${\mathcal{T}}_{q}$:
-$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}(q,{\mathcal{T}}_{q}).$ (14)
-After determining tables, for each $t\in\widehat{{\bm{T}}}_{q}$, let ${\mathcal{C}}^{t}_{q}$ represent the columns in table $t$. KaSLA then optimizes column linking within each selected table:
-$\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}(q,{\mathcal{C}}^{t}_{q}).$ (15)
+To tackle this optimization problem and achieve optimal linking results, we employ an efficient dynamic programming approach.
+The details of this algorithm, as implemented in KaSLA, are presented in Algorithm 1.
+Input: relevance score set ${\mathcal{R}}$, redundancy score set ${\mathcal{W}}$, the upper redundancy tolerance $u$
+Output: schema linking set $\hat{S}$
+$n\leftarrow|{\mathcal{R}}|$, $\hat{S}\leftarrow\{\}$
+[…]
+$keep\leftarrow\text{array of }(n+1)\times(u+1),\text{ initialized to False}$
+for $i\leftarrow 1$ to $n$ do
+       for $w\leftarrow 0$ to $u$ do
+             if ${\mathcal{W}}[i-1]\leq w$ then
+                   if $({\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]])>\mathcal{A}[i-1][w]$ then
+                         $\mathcal{A}[i][w]\leftarrow{\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]]$
+                         $keep[i][w]\leftarrow\text{True}$
+                  else
+                         $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
+            else
+                   $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
+$k\leftarrow u$  for $i\leftarrow n$ downto $1$ do
+       if $keep[i][k]$ then
+             $\hat{S}.\text{add}(i-1)$
+             $k\leftarrow k-{\mathcal{W}}[i-1]$
+return $\hat{S}$
+Algorithm 1 Dynamic Programming-based 0-1 Knapsack Optimization Algorithm
+IV-C2 Hierarchical Linking Strategy
+To ensure computational efficiency, KaSLA adopts a hierarchical schema linking strategy consisting of two stages: (1) table linking followed by (2) column linking within selected tables.
+This hierarchical approach significantly reduces the search space of candidate columns, enabling KaSLA to efficiently handle large-scale databases while maintaining robust performance.
+In the initial phase for a given query $q$ and the corresponding ${\mathcal{S}}_{q}$, KaSLA aims to identify the relevant tables from the set of all available tables in schema ${\mathcal{S}}_{q}$, denoted as ${\mathcal{T}}_{q}$:
+$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{T}}_{q}},{{\mathcal{W}}}_{{\mathcal{T}}_{q}},u^{{\mathcal{T}}}_{q}),$ (16)
+where ${{\mathcal{R}}}_{{\mathcal{T}}_{q}}$ and ${{\mathcal{W}}}_{{\mathcal{T}}_{q}}$ are the sets of estimated relevance and redundancy scores for tables in schema ${\mathcal{S}}_{q}$, and $u^{{\mathcal{T}}}_{q}$ is the upper redundancy tolerance for table linking.
+Different from the general schema elements ${\mathcal{S}}_{q}$ used in the equations in Section IV-B, the notation ${\mathcal{T}}_{q}$ and ${\mathcal{T}}$ specifically refer to tables in schema linking.
+After determining the selected tables, for each $t\in\widehat{{\bm{T}}}_{q}$, let ${\mathcal{C}}^{t}_{q}$ represent the columns in table $t$.
+KaSLA then optimizes column linking within each selected table:
+$\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}},{{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}},u^{{\mathcal{C}}}_{q}),$ (17)
+where ${{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}}$ and ${{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}}$ are the sets of estimated relevance and redundancy scores for columns in table $t$, and $u^{{\mathcal{C}}}_{q}$ is the upper redundancy tolerance for column linking.
 The final schema linking result $\widehat{{\bm{S}}}_{q}$ is the union of all selected tables and their corresponding selected columns, represented as:
 $\widehat{{\bm{S}}}_{q}=\widehat{{\bm{T}}}_{q}\cup\bigcup_{t\in\widehat{{\bm{T}}}_{q}}\widehat{{\bm{C}}}^{t}_{q}.$ (18)
-We provide a complexity analysis is detailed in Appendix D, offering insights into KaSLA’s implementation and computational efficiency.
-5 Experiments
+We present the complete procedure of the hierarchical linking strategy in Algorithm 2.
+Input: user query $q$, full schema ${\mathcal{S}}_{q}$
+Output: schema linking results $\widehat{{\bm{S}}}_{q}$, including a table linking set $\widehat{{\bm{T}}}_{q}$ and the corresponding column linking set $\widehat{{\bm{C}}}^{t}_{q}$ for each table $t\in\widehat{{\bm{T}}}_{q}$
+$({{\mathcal{R}}}_{{\mathcal{T}}_{q}},\left\{{{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}}\right\}_{t\in{\mathcal{T}}_{q}})=f_{\text{relevance}}(q,{\mathcal{S}}_{q})$,
+${{\mathcal{W}}}_{{\mathcal{T}}_{q}}=f_{\text{redundancy}}(q,{{\mathcal{R}}}_{{\mathcal{T}}_{q}}),(u^{{\mathcal{T}}}_{q},u^{{\mathcal{C}}}_{q})=f_{\text{tolerance}}(q)$
+$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{T}}_{q}},{{\mathcal{W}}}_{{\mathcal{T}}_{q}},u^{{\mathcal{T}}}_{q})$ // Table linking with Alg. 1
+for $t\in\widehat{{\bm{T}}}_{q}$ do
+       ${{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}}=f_{\text{redundancy}}(q,{{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}})$
+       $\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}},{{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}},u^{{\mathcal{C}}}_{q})$ // Column linking with Alg. 1
+return $\widehat{{\bm{S}}}_{q}=\widehat{{\bm{T}}}_{q}\cup\bigcup_{t\in\widehat{{\bm{T}}}_{q}}\widehat{{\bm{C}}}^{t}_{q}$
+Algorithm 2 Hierarchical Linking Strategy
+V Experiments
 We conducted comprehensive experiments to evaluate KaSLA and address the following research questions: RQ1: Does KaSLA enhance existing text-to-SQL baselines by improving their schema linking capabilities?
 RQ2: Does KaSLA demonstrate superior schema linking performance by handling element missing and redundancy?
-RQ3: Does each component in KaSLA contribute to the overall performance, and is KaSLA an efficient model for real-world text-to-SQL applications?
-5.1 Experiment Setup
-We conducted experiments on two well-know large-scale text-to-SQL datasets, BIRD (Li et al., 2023b) and Spider (Yu et al., 2018), incorporating multiple robust baselines based on LLMs with in-context learning and fine-tuning.
-For evaluating text-to-SQL, we utilized Execution Accuracy (EX) (Yu et al., 2018; Li et al., 2023b) and Valid Efficiency Score (VES) (Li et al., 2023b).
-We included multiple robust baselines from two categories:(i) LLMs + in-context learning baselines: C3-SQL (Dong et al., 2023), DIN-SQL (Pourreza and Rafiei, 2023), MAC-SQL (Wang et al., 2024), DAIL-SQL (Gao et al., 2024), SuperSQL (Li et al., 2024a), Dubo-SQL (Thorpe et al., 2024), TA-SQL (Qu et al., 2024), E-SQL (Caferoğlu and Ulusoy, 2024), CHESS (Talaei et al., 2024); and (ii) LLMs + fine-tuning baselines: DTS-SQL (Pourreza and Rafiei, 2024), CodeS (Li et al., 2024b).
-5.2 Main Results of SQL Generation (RQ1)
-To address RQ1, we conducted experiments to evaluate the SQL generation capabilities of state-of-the-art (SOTA) baselines enhanced by our KaSLA, with results presented in Table 1.
-KaSLA replaces the original schema linking component in models that already have schema linking capabilities or implements a full schema prompt substitution for those that do not.
-The results demonstrate that incorporating KaSLA consistently improves both Execution Accuracy (EX) and Valid Efficiency Score (VES) across baselines with diverse base models, including GPT-4, CodeS-15B, StarCoder2-15B, and Deepseek-v3.
-Notably, in the BIRD-dev dataset, KaSLA enhances the EX of CHESS with Deepseek-v3 from 61.34% to 63.30%.
-For local LLMs, both "CodeS + KaSLA" and "TA-SL + KaSLA" with CodeS-15B and StarCoder-15B achieve impressive increases in Execution Accuracy compared to their original versions.
-These results underscore KaSLA’s effectiveness and flexibility as a plug-in, showing that KaSLA plays an important role in improving complex text-to-SQL tasks.
-This adaptability not only confirms its robustness but also broadens its applicability across different models, achieving superior performance outcomes and making it a versatile addition to existing systems.
+RQ3: Does each component in KaSLA contribute to the overall performance, and what is the effect of parameter scaling?
+RQ4: Is KaSLA an efficient model for real-world text-to-SQL applications, especially in terms of its ability to transfer to unseen domains?
+V-A Experiment Setup
+V-A1 Datasets and Evaluation Metrics
+We conducted experiments on two well-known large-scale text-to-SQL datasets, BIRD [19] and Spider [20].
+Both datasets feature human-annotated queries and SQLs, complex database elements, and challenging cross-domain scenarios meticulously.
+For text-to-SQL evaluation, we measured performance using two well-established execution-based metrics: Execution Accuracy (EX) [20, 19] and Valid Efficiency Score (VES) [19].
+V-A2 Baselines
+We incorporating multiple robust baselines based on LLMs with in-context learning and fine-tuning.
+We included baselines from two categories:(i) LLMs + in-context learning baselines: C3-SQL [21], DIN-SQL [22], MAC-SQL [23], DAIL-SQL [4], SuperSQL [24], Dubo-SQL [25], TA-SQL [5], E-SQL [26], CHESS [11], RSL-SQL [12]; and (ii) LLMs + fine-tuning baselines: DTS-SQL [9], TA-SQL [5], CodeS [8].
+We also introduced Pure-SL, a baseline that directly fine-tunes LLMs for schema linking without any specially designed mechanisms.
+For baseline methods that include a schema linking module which can be cleanly separated from their main framework, we distinguish these variants by adding an ’-SL’ suffix to their original names.
+For example, a baseline named “CodeS” would become “CodeS-SL” when referring specifically to its schema linking component.
+V-A3 Implementation Details
+We implemented the proposed KaSLA and baselines under the following settings: (i) Close-source LLM-based experiments: We conducted experiments for most in-context learning baselines using GPT-4-turbo.
+For E-SQL, CHESS, and RSL-SQL, which originally used GPT-4o, we switched to deepseek-V3 because the API cost of GPT-4o was too high for these methods.
+Since previous work has shown that deepseek-V3 achieves similar performance to GPT-4o at a much lower cost [27].
+(ii) Open-source LLM-based experiments: We use CodeS-15B [8] and StarCoder2-15B [28] as the base LLMs for the SQL generation model, and reproduce all baselines using LoRA [15] for parameter-efficient fine-tuning.
+The learning rate is set to $41e-$ 4 with cosine decay, the batch size is 16, and the training runs for 3 epochs.
+(iii) KaSLA: We set the top-$K$ value for redundancy tolerance estimation to 30.
+For the scoring models in KaSLA, we fine-tune a RoBERTa-Large [16] as the binary scoring model following [8], and a Deepseek-coder-1.3B [14] as the probabilistic scoring model using the same fine-tuning settings with all fine-tuning baselines as described in (ii).
+TABLE I: The SQL generation performance of enhanced text-to-SQL models using KaSLA, which features 1.6 B parameters combining DeepSeek-coder-1.3B and RoBERTa-Large, is evaluated in terms of Execution Accuracy (EX) (%) and Valid Efficiency Score (VES) (%) on the BIRD-dev and Spider-dev datasets.
+We use KaSLA to replace the original schema linking component or the full schema in the baseline models, denoted as “+ KASLA”, and the relative improvements are highlighted with a gray background.
 Base Model Method BIRD-dev Spider-dev
@@
 CodeS-15B CodeS 65.19 48.60 38.19 57.63 63.22 94.76 91.26 72.99 64.46 84.72 83.52
-CodeS + KaSLA 68.32 50.75 38.19 60.17 64.52 96.37 89.46 76.44 63.86 84.82 84.46
-StarCoder2 -15B TA-SL 59.89 45.38 34.72 53.13 60.51 95.56 92.38 78.16 63.25 85.98 85.03
-TA-SL + KaSLA 66.59 49.46 38.89 58.80 63.56 94.76 92.15 77.59 66.87 86.27 85.47
+CodeS + KaSLA 68.32 50.75 38.19 60.17 $\uparrow$ 4.41% 64.52 $\uparrow$ 2.06% 96.37 89.46 76.44 63.86 84.82 $\uparrow$ 0.12% 84.46 $\uparrow$ 1.13%
+StarCoder2 -15B TA-SQL 59.89 45.38 34.72 53.13 60.51 95.56 92.38 78.16 63.25 85.98 85.03
+TA-SQL + KaSLA 66.59 49.46 38.89 58.80 $\uparrow$ 10.67% 63.56 $\uparrow$ 5.04% 94.76 92.15 77.59 66.87 86.27 $\uparrow$ 0.34% 85.47 $\uparrow$ 0.52%
 DTS-SQL 63.03 46.02 34.72 55.22 64.17 91.94 90.58 78.74 66.87 85.11 85.49
-DTS-SQL + KaSLA 63.89 50.11 39.58 57.43 65.20 94.76 90.81 80.46 68.07 86.36 85.81
-CodeS 68.00 51.40 39.58 60.30 65.04 94.76 91.26 72.99 64.46 84.72 83.52
-CodeS + KaSLA 67.68 54.84 42.36 61.41 66.87 94.76 91.03 78.74 66.27 85.88 84.46
+DTS-SQL + KaSLA 63.89 50.11 39.58 57.43 $\uparrow$ 4.00% 65.20 $\uparrow$ 1.61% 94.76 90.81 80.46 68.07 86.36 $\uparrow$ 1.47% 85.81 $\uparrow$ 0.37%
 Deepseek -V3 E-SQL 67.24 51.61 41.67 60.10 62.16 95.57 91.93 72.99 64.46 85.21 83.81
-E-SQL + KaSLA 69.40 53.55 43.75 62.18 65.32 95.97 92.38 73.56 65.06 85.69 85.72
+E-SQL + KaSLA 69.40 53.55 43.75 62.18 $\uparrow$ 3.46% 65.32 $\uparrow$ 5.08% 95.97 92.38 73.56 65.06 85.69 $\uparrow$ 0.56% 85.72 $\uparrow$ 2.28%
 CHESS 67.89 52.90 46.53 61.34 65.28 95.57 92.16 73.56 65.06 85.50 85.33
-CHESS + KaSLA 69.94 54.84 47.92 63.30 67.21 96.37 92.83 74.14 65.66 86.17 86.35
-Table 1: The SQL generation performance of enhanced text-to-SQL models using KaSLA, which features 1.6 B parameters combining DeepSeek-coder-1.3B and RoBERTa-Large, is evaluated in terms of Execution Accuracy (EX) (%) and Valid Efficiency Score (VES) (%) on the BIRD-dev and Spider-dev datasets.
-We utilize KaSLA to substitute the original schema linking component in models that already incorporate schema linking, or to replace the full schema prompt for models that do not.
-5.3 Experimental Results of Schema Linking (RQ2)
-To answer the RQ2, we conduct the experiments and report the evaluation results of schema linking methods on BIRD-dev in Figure 2.
-We measured the performance of these schema linking methods in both table linking and column linking.
-The results indicate that, under the enhanced F1 score which specifically evaluates missing elements, only KaSLA achieves high accuracy in the challenging task of column linking.
-Schema linking models using LLMs often perform well in table linking but fall short in column linking due to the lack of additional optimizations for missing elements.
-This highlights the advantages of KaSLA’s framework and demonstrates how enhanced metrics can reveal the significant issue of element missing in current schema linking practices.
-Schema linking method Table linking Column linking
-DTS-SL (starcoder2-15b) 61.94 23.71
-CodeS-SL (starcoder2-15b) 58.21 29.56
-RSL-SQL (deepseek-v3) 67.11 17.20
-KaSLA (deepseek-coder-1.3B) 81.92 62.78
-Table 2: Schema linking performance with the proposed enhanced linking metric on BIRD-dev.
-Ablation EX
-CodeS + KaSLA 60.17
+CHESS + KaSLA 69.94 54.84 47.92 63.30 $\uparrow$ 3.20% 67.21 $\uparrow$ 2.96% 96.37 92.83 74.14 65.66 86.17 $\uparrow$ 0.78% 86.35 $\uparrow$ 1.20%
+RSL-SQL 69.95 54.62 47.22 63.17 67.11 96.37 93.05 74.71 66.87 86.56 86.42
+RSL-SQL + KaSLA 71.03 56.34 51.39 64.73 $\uparrow$ 2.47% 68.12 $\uparrow$ 1.35% 97.18 93.50 75.29 68.07 87.14 $\uparrow$ 1.50% 87.36 $\uparrow$ 1.09%
+V-B Main Results of SQL Generation (RQ1)
+To address RQ1, we conducted experiments to evaluate the SQL generation capabilities of state-of-the-art (SOTA) baselines enhanced by our KaSLA, with results presented in Table I.
+KaSLA serves as a plug-in method: for models with built-in schema linking modules, it replaces their original linking component; for models without a schema linking component, it substitutes the conventional full schema prompting approach.
+TABLE II: Schema linking performance with Recall+, Precision+, and F1+ on BIRD-dev.
+We extracted the schema linking components from the baselines and used “-SL” instead of the original “-SQL” to indicate this.
+Dataset Metric Recall-based Method Generative Method KaSLA
+CodeS-SL ICL DIN-SL DTS-SL CHESS-SL E-SL Pure-SL TA-SL RSL-SL
+BIRD-dev Recall+ Table 99.15 88.40 82.66 64.99 80.70 69.04 64.99 79.47 92.57 94.52
+Column 85.14 22.03 16.17 62.52 14.93 22.36 41.33 54.69 57.11 81.29
+Precision+ Table 43.03 82.75 78.80 60.29 49.92 68.01 60.29 74.49 83.89 75.00
+Column 18.65 20.69 15.12 16.50 7.09 21.57 37.23 52.31 35.14 53.53
+F1+ Table 58.21 84.79 80.18 61.94 58.51 68.39 61.94 76.16 86.93 81.92
+Column 29.56 21.20 15.51 23.71 9.03 21.87 38.77 53.20 40.60 62.78
+The results in Table I demonstrate that incorporating KaSLA consistently improves both Execution Accuracy (EX) and Valid Efficiency Score (VES) across baselines with diverse base models, including GPT-4, CodeS-15B, StarCoder2-15B, and Deepseek-v3.
+Notably, on the BIRD-dev dataset, KaSLA raises the EX of RSL-SQL with Deepseek-v3 from 63.17% to 64.73%, achieving a 2.47% improvement and the best overall result.
+KaSLA also delivers a 3.20% increase on CHESS.
+For local LLMs, both “CodeS + KaSLA” with CodeS-15B and “TA-SL + KaSLA” with StarCoder-15B achieve significant gains in Execution Accuracy compared to their original versions.
+We also found that the differences between baselines on Spider-dev are smaller compared to BIRD-dev.
+This may be because the schemas in Spider-dev are less complex.
+Even in this case, KaSLA can still consistently provide improvements.
+These results underscore KaSLA’s effectiveness and flexibility as a plug-in, showing that KaSLA plays an important role in improving complex text-to-SQL tasks.
+This adaptability not only confirms its robustness but also broadens its applicability across different models, achieving superior performance outcomes and making it a versatile addition to existing text-to-SQL systems.
+V-C Experimental Results of Schema Linking (RQ2)
+To answer the RQ2, we conduct the evaluation of schema linking methods on BIRD-dev using the proposed schema linking metrics.
+Table II shows the performance of various schema linking methods on BIRD-dev, evaluated by Recall+, Precision+ and F1+ at both the table and column levels.
+The results allow us to compare the schema linking performance of recall-based, generative, and the proposed KaSLA method.
+As shown in Table II, most current schema linking models perform well in table linking, but linking relevant columns from a large number of candidate columns remains challenging.
+The recall-based method (CodeS-SL) achieves the highest Recall+ at both the table and column levels, indicating it is very effective at identifying relevant elements and avoiding missing important ones.
+However, this comes at the cost of low Precision+, especially for columns, meaning many redundant elements are also linked.
+As a result, its F1+ scores for both tables and columns are low.
+Generative methods tend to optimize for Precision+, but their Recall+ is much lower than that of the recall-based method, indicating more missing elements.
+Even the best-performing generative model, RSL-SL, achieves a low Precision+ and, therefore, a low F1+ on the challenging column linking task.
+Notably, models like E-SQL and CHESS, which achieve strong results in Table I, obtain low scores for column linking when their schema linking modules are evaluated separately.
+This suggests that E-SQL uses LLMs to process the full schema directly without schema linking [26] and CHESS reduces error rates through multiple rounds of iteration [11], which both result in high LLM API costs.
+KaSLA demonstrates strong and balanced performance.
+Compared to generative models, it achieves higher Recall+, only slightly lower than the recall-based method, while also obtaining competitive Precision+ and the best F1+ at the column level, as well as a strong table-level F1+.
+These results show that KaSLA achieves the best overall trade-off by effectively balancing recall and precision, successfully reducing missing and redundant elements simultaneously.
+This allows KaSLA to outperform recall-based and generative baselines in schema linking quality, particularly for the challenging task of column linking, thus demonstrating the advantage of its knapsack-optimization framework for text-to-SQL tasks.
+V-D Ablation Study (RQ3)
+To answer RQ3, we present: (i) the ablation study results for KaSLA’s key components, (ii) the performance analysis with different language model choices for the probabilistic scoring model, and (iii) the effects of scaling up the scoring model.
+V-D1 Key Components in KaSLA
+As shown in Table III: (i) Removing the binary scoring model results in a lack of confirmation for high-confidence elements, leading to the inclusion of more redundant ones.
+This underscores the binary model’s role in filtering out less relevant elements by ensuring that elements with high certainty are correctly identified.
+(ii) Omitting the probabilistic scoring model risks missing relevant elements, because the soft relevance probabilities it provides can effectively protect against the mistakes made by the binary scoring model with hard relevance scores.
+It provides a more robust estimation.
+(iii) The hierarchical strategy effectively reduces the number of candidate columns.
+Without this strategy, column linking would have to be performed on all candidate columns, which introduces extra noise and increases computational cost.
+(iv) Without the upper limit set at 1, the model may introduce disproportionate comparisons among relevant items, potentially excluding some relevant ones.
+This constraint is crucial to maintain balanced relevance estimation, preventing the overshadowing of elements that might otherwise contribute positively to SQL generation.
+These ablation study highlights the significance of key components within KaSLA.
+TABLE III: Ablation studies of the key components of KaSLA. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS-15B as the text-to-SQL backbone.
+Dataset Ablation of key components of KaSLA EX
+BIRD-dev CodeS + KaSLA 60.17
 w/o Binary scoring model 56.06
@@
 w/o upper limit 1 in Eq. 12 53.26
-Table 3: Ablation studies of the key components of KaSLA. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
-Dataset Schema linking model Binary scoring function with deepseek-coder-1.3B Probabilistic scoring model with RoBERTa-Large Estimation and dynamic programming Text-to-SQL Total time EX
-BIRD-dev Full Schema / / / 5.15 s 5.15 s 57.63
-KaSLA 3.5 s 0.12 s $<$ 0.01 s 1.85 s 5.48 s 60.17
-Table 4: Inference time cost per instance of each component in KaSLA using CodeS-15B as the text-to-SQL model.
-Model SuperSQL CodeS CodeS + KaSLA CodeS + KaSLA (Transfer)
-BIRD-dev 58.60 60.30 60.17 60.02
+(a) Binary scoring model
+(b) Probabilistic scoring model
+Figure 6: (a) Execution accuracy (EX, %) of CodeS + KaSLA on BIRD-dev with different language models as the binary scoring model.
+(b) Effect of model parameter scaling of probabilistic scoring model with Deepseek-coder.
+All experiments use CodeS-15B as the text-to-SQL backbone.
+V-D2 Binary Scoring Model
+We present ablation studies on the choice of language models for the binary scoring model used in KaSLA, as detailed in Figure 6(a).
+We evaluate the SQL generation performance of CodeS-15B on BIRD-dev using KaSLA with different language models as the binary scoring model.
+In Figure 6(a), we find that RoBERTa-Large achieves the best results.
+One reason is that RoBERTa enhances BERT by employing a more comprehensive training process, leading to improved text representation abilities [16].
+This allows RoBERTa-Large to provide better generalization and understanding of language nuances, which are crucial for schema linking tasks where capturing subtle semantic relationships is important.
+Moreover, reducing the number of parameters (e.g., using BERT-base or RoBERTa-base) resulted in relatively underfitted models that failed to capture sufficient semantic information for the task.
+Conversely, increasing the number of parameters (e.g., SGPT-1.3B) introduced a tendency toward overfitting, which did not yield performance improvements.
+These results indicate that RoBERTa-Large offers an effective balance between model capacity and generalization, making it an excellent choice for the binary scoring model.
+TABLE IV: The comparison of schema linking performance between CodeS-SL and KaSLA is shown using a randomly selected case from BIRD-dev.
+We highlight the missing relevant element in CodeS-SL’s schema linking results in blue, and the redundant elements that mislead the SQL generation model in red.
+These two types of errors together result in the incorrect SQL generation of CodeS-SL.
+In contrast, KaSLA successfully avoids these issues.
+Question ‘Posters’ is the expense description; ‘September Speaker’ is an event name; Indicate the cost of posters for ‘September Speaker’ event.
+Method CodeS-SL KaSLA
+Schema linking result event: [event_id, event_name, event_date, status, notes, location] zip_code: [type, county, city, state, zip_code, short_state] expense: [cost, expense_description, expense_id, expense_date, approved, link_to_budget] attendance: [link_to_member, link_to_event] budget: [event_status, category, spent, amount, remaining, link_to_event] event: [event_id, event_name] expense: [cost, expense_description, expense_date, link_to_budget] budget: [budget_id, link_to_event]
+Missing elements budget: [budget_id] \
+Redundant elements event: [event_date, status, notes, location] zip_code: [type, county, city, state, zip_code, short_state] expense: [expense_id, expense_date, approved] attendance: [link_to_member, link_to_event] budget: [event_status, category, spent, amount, remaining] expense: [expense_date]
+Generated SQL using CodeS-15B SELECT expense.cost FROM expense INNER JOIN attendance ON expense.link_to_member = attendance.link_to_member INNER JOIN event ON attendance.link_to_event = event.event_id WHERE event.event_name = ‘September Speaker’ AND expense.expense_description = ‘Posters’ SELECT expense.cost FROM event INNER JOIN budget ON event.event_id = budget.link_to_event INNER JOIN expense ON budget.budget_id = expense.link_to_budget WHERE event.event_name = ‘September Speaker’ AND expense.expense_description = ‘Posters’
+SQL execution result Incorrect Correct
+V-D3 Probabilistic Scoring Model
+We present ablation studies on LLM selection with parameter scaling for KaSLA’s probabilistic scoring component in Figure 6(b).
+We find that although using larger LLMs as the probabilistic scoring model leads to improved performance, the difference in execution accuracy between Deepseek-coder-1.3B and the larger 6.5B and 7B models is marginal.
+Importantly, Deepseek-coder-1.3B offers clear advantages in terms of lower resource consumption and faster inference time due to its smaller parameter size.
+These results indicate that, within the knapsack optimization framework of KaSLA, a low-parameter LLM can achieve performance comparable to that of much larger models.
+This finding demonstrates that KaSLA is well-suited for efficient and scalable deployment in practical text-to-SQL applications, where smaller models are often preferred due to constraints on computational resources and the need for faster inference.
+V-E Empirical Studies (RQ4)
+To address RQ4, we conduct a series of empirical studies to evaluate the practical effectiveness of KaSLA.
+First, we provide a randomly selected case study in Table IV to demonstrate how KaSLA improves SQL generation in real scenarios compared with a baseline.
+Next, we evaluate KaSLA’s transferability across domains and scenarios, with the results summarized in Table V and Table VI, respectively.
+The results show that KaSLA has strong transferability on unseen data.
+We also offer a complexity analysis and present the inference time cost when integrating KaSLA into text-to-SQL frameworks in Table VII, highlighting KaSLA’s ability to enhance SQL generation accuracy with only a modest increase in inference time.
+Finally, we systematically analyze the failure cases of “CodeS + KaSLA” on the BIRD-dev, as illustrated in Figure 8, to provide a comprehensive understanding of KaSLA’s schema linking performance in SQL generation tasks.
+V-E1 Case Study
+To illustrate KaSLA’s practical impact, we provide a randomly selected case study ib BIRD-dev dataset to illustrate how KaSLA improves SQL generation in practice, comparing with CodeS-SL.
+The case are detailed in Table IV, it demonstrates how KaSLA effectively avoids missing relevant elements compared to the baseline.
+The original schema linking model in CodeS introduced many redundant elements while missing the relevant column $budget\_id$ in the $budget$ table.
+After replacing it with KaSLA, the schema linking results included all relevant elements with minimal redundancy, enabling the SQL generation model to produce the correct query.
+This case highlights KaSLA’s ability to handle complex, real-world scenarios, which complements the quantitative results with tangible evidence of its utility.
+Figure 7: The seen and unseen data in BIRD-dev relative to BIRD-train.
+TABLE V: Execution Accuracy (EX) (%) of text-to-SQL models evaluated on seen data and un-seen data in BIRD-dev.
+Evaluation dataset SuperSQL CodeS CodeS + KaSLA
+Original BIRD-dev 58.60 57.63 60.17
+Seen data in BIRD-dev 69.90 67.58 71.06
+Un-seen data in BIRD-dev 49.35 51.01 86.48
+V-E2 Transferability
+In real-world scenarios, databases come from diverse domains, leading to varied data and queries.
+Unlike open-source datasets, real-world tasks often lack enough in-domain training data, so cross-domain transferability of schema linking models is essential.
+We specifically evaluate KaSLA’s performance in cross-domain and cross-scenario settings.
+Cross-domain Transferability
+The BIRD dataset is specifically designed for cross-domain evaluation.
+As shown in Figure 7, there are 5 databases in BIRD-dev whose domains match those in BIRD-train.
+For example, “european_football_2” in BIRD-dev matches with “european_football_1” in BIRD-train.
+More relations between data from BIRD-dev and BIRD-train are illustrated in Figure 7.
+This type of data is considered seen data, and the model can learn prior knowledge for these domains from BIRD-train.
+In contrast, there are six databases in BIRD-dev that do not have directly related domains in BIRD-train.
+For example, “card_games” in BIRD-dev does not have any domain-matched database in BIRD-train.
+For these databases, the model cannot acquire relevant prior knowledge through training.
+Evaluating KaSLA on these unseen domains provides a actual test of its cross-domain transferability.
+We evaluated the SQL generation performance of KaSLA and baseline models on data from both the unseen and seen databases in BIRD-dev.
+As shown in Table V, KaSLA demonstrates a clear advantage over baselines on both seen and unseen data.
+This strong performance is attributed to the robustness of the KaSLA optimization framework, which uses knapsack optimization to reduce missing and redundant elements.
+It enables KaSLA to achieve optimal results even on unseen domains after training on open-source datasets.
+Cross-scenario Transferability
+Given the diversity of databases in real-world scenarios, evaluating KaSLA’s performance under cross-scenario conditions provides a more practical assessment of its potential for real-world text-to-SQL applications.
+Since BIRD and Spider are datasets from different scenarios, we assess KaSLA’s cross-scenario transferability by evaluating KaSLA trained on BIRD-train with Spider-dev, and vice versa—evaluating KaSLA trained on Spider-train with BIRD-dev.
+For baselines, we follow the standard setting of using training or prompt sampling from the same scenario as the evaluation data.
+Results are shown in Table VI.
+Here, “KaSLA (Transfer)” denotes KaSLA trained on the training set of a scenario different from the evaluation data.
+We observe that KaSLA trained on a different-scenario dataset still outperforms the baselines and only slightly underperforms compared to KaSLA trained with domain-specific data from the matching scenario.
+These results demonstrate that KaSLA has strong cross-scenario transferability, highlighting its robustness and adaptability when applied to new text-to-SQL scenarios.
+This also confirms that KaSLA, when trained on open-source datasets, can be directly and effectively applied to unseen scenarios in real-world applications.
+TABLE VI: Execution Accuracy (EX) (%) of KaSLA trained on Spider-train but evaluated on BIRD-dev, and vice versa, for cross-scenario transfer ability evaluation.
+Evaluation dataset SuperSQL CodeS CodeS + KaSLA CodeS + KaSLA (Transfer)
+BIRD-dev 58.60 57.63 60.17 60.02
 Spider-dev 87.40 84.72 84.82 84.56
-Table 5: Execution Accuracy (EX) (%) of KaSLA trained on Spider-train but evaluated on BIRD-dev, and vice versa, for cross-scenario transfer ability evaluation.
-5.4 Empirical studies (RQ3)
-Ablation Study.
-We presented the ablation study results for the key components in KaSLA in Table 3, the results concerning the choice of language model for the probabilistic scoring model are shown in Table 8 in Apendix and the scaling up performance of the scoring model are shown in Table 7 in Apendix, respectively.
-As shown in Table 3: (i) Removing the binary scoring model results in a lack of confirmation for high-confidence elements, leading to the inclusion of more redundant ones.
-This underscores the binary model’s role in filtering out less relevant elements by ensuring that elements with high certainty are correctly identified.
-(ii) Omitting the probabilistic scoring model risks missing potentially relevant elements, resulting in gaps within schema linking.
-The probabilistic model is essential for assigning a non-zero probability to all elements, thus capturing those with potential relevance.
-(iii) The hierarchical strategy effectively reduces the number of candidate columns, thereby lowering complexity and improving efficiency.
-By structuring the linking process in stages, KaSLA can more precisely target relevant schema components.
-(iv) Without the upper limit set at 1, the model introduces disproportionate comparisons among similarly relevant items, potentially excluding some relevant elements.
-This constraint is crucial to maintain balanced relevance assessments, preventing the overshadowing of items that might otherwise contribute positively to SQL generation.
-These ablation study highlights the significance of key components within KaSLA.
-Transferability.
-We conducted two experiments to evaluate KaSLA’s performance in cross-scenario transfer.
-We trained the binary and probabilistic scoring models on the Spider training dataset and evaluated it on the BIRD dev dataset, and vice versa, to explore its cross-scenario transfer ability.
-Results are provided in Table 5.
-We can find that pre-training KaSLA on public datasets yields results that outperform the baselines and are only slightly lower than what domain-specific fine-tuning would achieve.
-The results show that KaSLA demonstrates strong cross-scenario transferability, highlighting its robustness and adaptability across different data domains.
-Inference time cost.
-We evaluated the inference time cost of integrating KaSLA into text-to-SQL frameworks in Table 4.
+V-E3 Complexity Analysis and Inference Time Cost
+We provide a complexity analysis for offering insights into KaSLA’s implementation and computational efficiency.
+The efficiency of KaSLA stems from its hierarchical approach, dynamic programming and inherent parallelism.
+KaSLA’s time complexity is determined by two main phases: table linking and column linking within the linked tables.
+For a database with $|{\mathcal{T}}|$ tables and a table tolerance of $u^{{\mathcal{T}}}$, the table linking phase has a time complexity of $O(|{\mathcal{T}}|u^{{\mathcal{T}}})$.
+KaSLA then links columns within the selected table.
+Let $|{\mathcal{C}}^{t}|$ and $u^{{\mathcal{C}}}$ represent the number of columns and the column capacity of a selected table, respectively.
+The time complexity of the column linking phase is $O(|{\mathcal{C}}^{t}|u^{{\mathcal{C}}})$.
+Consequently, the total time complexity of KaSLA is $O(|{\mathcal{T}}|u^{{\mathcal{T}}}+|{\mathcal{T}}|(|{\mathcal{C}}^{t}|u^{{\mathcal{C}}}))$.
+Since all factors are constants, this results in linear complexity.
+We evaluated the inference time cost of integrating KaSLA into text-to-SQL frameworks in Table VII.
 By utilizing the lightweight LLM DeepSeek-coder-1.3B and significantly reducing the number of tokens in the input prompts through optimized schema linking, KaSLA achieves improved SQL generation accuracy with only a modest increase in inference time.
 This showcases KaSLA’s efficiency in balancing performance and computational demands.
-6 Related Work
-Schema linking is essential for aligning natural language queries with database components, playing a pivotal role in the overall text-to-SQL process.
-Recent approaches utilize large language models (LLMs) to enhance schema linkage through methods such as semantic retrieval, historical query integration, and bidirectional linking (Pourreza and Rafiei, 2024; Cao et al., 2024; Zhang et al., 2024).
-Techniques like DAIL-SQL and E-SQL refine linkage by leveraging LLMs’ contextual understanding and query enrichment capabilities (Gao et al., 2024; Caferoğlu and Ulusoy, 2024).
-Retrieval-based strategies focus on retrieve information from database to improve SQL generation (Pourreza and Rafiei, 2024).
-Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) utilize semantic retrieval to improve schema linking by accessing relevant database information.
+TABLE VII: Inference time cost per instance of each component in KaSLA using CodeS-15B as the text-to-SQL model.
+Model Full Schema KaSLA
+Binary scoring function / 3.5 s
+Probabilistic scoring model / 0.12 s
+Estimation and dynamic programming / $<$ 0.01 s
+SQL generation with CodeS-15B 5.15 s 1.85 s
+Total time 5.15 s 5.48 s
+EX 57.63 60.17
+V-E4 Failure Case Analysis
+Figure 8: Failure case analysis of “CodeS + KaSLA” in BRD-dev.
+We present a systematic analysis of the failure cases of “CodeS + KaSLA” in BRD-dev, as detailed in Figure 8, to comprehensively study the schema linking performance of KaSLA.
+We first categorize all evaluation results into two groups: correct SQL generation and incorrect SQL generation.
+Next, we further divide the cases of incorrect SQL generation into three subcategories: (a) schema linking performs well, but the SQL generation model is insufficient; (b) the SQL generation model is misled by redundant elements; and (c) the SQL generation model is misled by missing elements.
+Notably, subcategories (b) and (c) are directly caused by errors in the schema linking model, which subsequently lead to incorrect SQL generation.
+In contrast, in category (a), the schema linking provides correct linking results, yet the SQL generation model still produces incorrect SQL.
+As show in Figure 8, in 79.40 % of examples (60.17% + 19.23%), KaSLA provides schema links without missing or misleading redundant elements.
+However, while 60.17% of these cases yield correct SQL, 19.23% still result in incorrect SQL due to limitations in the SQL generation model.
+This suggests that adopting more advanced SQL generation models could further improve the text-to-SQL performance.
+We can also observe that, despite KaSLA’s tailored optimization for schema linking, it still misses relevant schema elements in 9.52% of examples and introduces redundant elements, which mislead the SQL generation model in 11.08% of cases.
+This indicates room for improvement in the future work, particularly in refining the knapsack optimization algorithm or enhancing the relevance scoring model’s accuracy to reduce redundancies while minimizing the omission of relevant elements.
+VI Related Work
+VI-A LLM-based Text-to-SQL
+Alongside the emergence of LLMs, in-context learning (ICL), also known as prompt engineering has become a core method driving advancements [29].
+Early efforts in LLM-based text-to-SQL studied the effectiveness of ICL with different prompt designs on various LLMs’ structures [30].
+With the success of well-designed ICL methods in other natural language understanding tasks [31, 32], their integration into text-to-SQL has also achieved solid improvements [33].
+Chain-of-Thought (CoT) prompting [31] shows great potential for text-to-SQL tasks: DIN-SQL [22] and ACT-SQL [34] proposed a CoT-based framework to enhance SQL generation.
+To optimize the prompt, DAIL-SQL [4] and Knowledge-to-SQL [35] provide related samples and helpful knowledge for LLMs to learn with.
+Furthermore, MAC-SQL [23] and CHESS [11] employs a multi-agent collaboration framework to generate and refine SQL; CHASE-SQL [36] proposes an ICL prompt for question decomposition and involves a novel online synthetic example generation.
+ICL-based methods are the mainstream paradigm of LLM-based text-to-SQL, which have made significant progress and continue to be widely studied in the most current work [3].
+As another paradigm, fine-tuning is an intuitive and widely recognized technique for enabling LLMs to perform specific downstream tasks [32].
+Even though code-specific LLMs are trained on extensive programming scenarios, they still struggle when facing challenging user queries and complex database environments [37, 38, 19].
+As a straightforward method, supervised fine-tuning (SFT) [32] is utilized to adapt open-source LLMs for SQL generation [4].
+For well-designed methods, CodeS [8] proposes a two-stage training framework achieving impressive performance with open-source LLMs; PURPLE [39] retrieves demonstrations containing the requisite logical operator composition to guide LLMs for SQL generation.
+DTS-SQL [9] fine-tunes two LLMs separately for schema linking generation and SQL generation; XiYan-SQL [40] proposes a multi-generator ensemble framework through fine-tuning to enhance SQL candidate quality and diversity, with a refined selection model and schema representation.
+VI-B Schema Linking in Text-to-SQL
+Within the text-to-SQL task, schema linking is proven as an essential module for aligning natural language queries with database contents, playing a pivotal role in the overall SQL generation process [3].
+Recent approaches utilize LLMs to enhance schema linkage through methods such as semantic retrieval, historical query integration, and bidirectional linking [9, 12, 41].
+Techniques like DAIL-SQL [4] and E-SQL [26] refine linkage by leveraging LLMs’ contextual understanding and query enrichment capabilities [4, 26].
+Retrieval-based strategies focus on retrieving information from the database to improve SQL generation [9].
+CodeS [8] utilizes semantic retrieval to improve schema linking by accessing relevant database information.
+CHESS [11] design dedicated schema selector agents that prompt LLMs to choose relevant tables and columns.
 Despite these advancements, many state-of-the-art text-to-SQL models emphasize fine-tuning strategy.
-DTS-SQL (Pourreza and Rafiei, 2024) fine-tunes a local LLM to select relevant tables from user queries, while Solid-SQL (Liu et al., 2025) introduces a data augmentation method for enhancing schema linking.
-DELLM (Hong et al., 2024a) finetuned a specialized data expert LLM to provide additional knowledge for schema linking and SQL generation.
-Approaches like TA-SQL (Qu et al., 2024) use in-context learning to guide LLMs, generating dummy SQL for linked element abstraction.
-Similarly, RSL-SQL (Cao et al., 2024) employs a bidirectional approach for schema simplification.
-Meanwhile, E-SQL (Qu et al., 2024) and (Maamari et al., 2024) note that schema linking challenges persist, leading some models to use the entire schema directly as input for SQL generation.
+DTS-SQL [9] fine-tunes a local LLM to select relevant tables from user queries, while Solid-SQL [10] introduces a data augmentation method for enhancing schema linking.
+Approaches like TA-SQL [5] use in-context learning to guide LLMs, generating dummy SQL for linked element abstraction.
+Similarly, RSL-SQL [12] employs a bidirectional approach for schema simplification.
+Meanwhile, E-SQL [5] and Distillery-SQL [6] note that schema linking challenges persist, leading some models to use the entire schema directly as input for SQL generation.
 Different with the previous research, our KaSLA utilize a knapsack optimization framework to avoid the relevant elements missing and the inclusion of redundant elements, optimizing schema linking to enhance SQL generation accuracy effectively.
-7 Conclusion
-This paper introduced the Knapsack Schema Linking Agent (KaSLA), a pioneering approach to overcoming schema linking challenges in text-to-SQL tasks.
-Based on knapsack optimization, KaSLA effectively prevent the missing of relevant elements and excess redundancy, thereby significantly enhancing SQL generation accuracy.
-Our introduction of an enhanced schema linking metric, sets a new standard for evaluating schema linking performance.
+VII Conclusion
+This paper studies the schema linking in text-to-SQL, specifically how to link relevant elements while avoiding missing and redundant ones.
+We first proposed Recall+ and Precision+ with a restricted missing-element indicator to better assess linking performance.
+Furthermore, we introduced the Knapsack Schema Linking Approach (KaSLA), a pioneering approach to overcoming schema linking challenges in text-to-SQL tasks.
+Based on knapsack optimization, KaSLA effectively links relevant elements while preventing missing and redundant elements simultaneously, thereby significantly improving SQL generation accuracy.
 KaSLA utilizes a hierarchical linking strategy, starting with optimal table linking and then proceeding to column linking within selected tables, effectively reducing the candidate search space.
@@
 These findings underscore KaSLA’s potential to revolutionize schema linking and advance the broader capabilities of text-to-SQL systems, offering a robust solution that paves the way for more precise and efficient database interactions.
-8 Limitation
-There are mainly two limitations of this work.
-First, although KaSLA demonstrates significant advancements in schema linking accuracy and efficiency, its performance traditionally depends on a comprehensive training dataset with detailed ground truth linking results.
-However, our experiments indicate that KaSLA possesses a degree of transferability, showing promise even in scenarios where training data is sparse or less diverse.
-This adaptability suggests that KaSLA can maintain reasonable effectiveness across different database environments, although performance might still be impacted in highly dynamic or rapidly evolving schema structures.
-Additionally, the reliance on a fine-tuning process of LLMs, integrating DeepSeek-coder-1.3B, could pose scalability challenges in resource-constrained settings.
-Despite these limitations, KaSLA remains a powerful tool for schema linking optimization in text-to-SQL tasks, and future work could focus on enhancing its adaptability and efficiency.
-9 Ethics Statement
-We confirm that we have fully complied with the ACL Ethics Policy in this study. All the datasets are publicly available and have been extensively used in research related to text-to-SQL.
-Appendix A Implementation Details.
-We implemented the proposed KaSLA and baselines under the following settings: For top-$K$ used in the redundancy tolerance estimation of KaSLA, we set it to 30.
-For method using fine-tuned LLM, We utilized LoRA (Hu et al., 2021) for parameter-efficient fine-tuning.
-The learning rate was initialized to $41e-$ 4 with a cosine decay, the batch size was set to 16, and the training epoch was 3.
-Appendix B Input formats
-We provided the input and output formats of the binary scoring function of KaSLA introduced in section 4.1 in Table 6.
-Type Input format
-Full schema table department, columns = [ department__\__id (int |||| primary key||||values: 1), name (text||||values: State, Treasury), ranking (int |||| values: 1, 2)] table head, columns = [ head__\__id (int |||| primary key |||| values: 1), name (text |||| values: Tiger Woods), age (real |||| values: 67.0)] table management, columns = [ department__\__id ( int |||| primary key |||| values: 2), head__\__id (int |||| primary key |||| values: 5), temporary__\__acting (text |||| values: Yes)] foreign keys: management.head__\__id = head.head__\__id, management.department__\__id = department.department__\__id
-Question How many heads of the departments are older than 56? Generate SQL to solve the above question and list the relevant columns in each table
-Type Output format
-Simulated SQL SELECT count(*) FROM head WHERE age > 56
-Schema linking results "department": null, "management": null, "head": [ "age" ]
-Table 6: Input and output formats of the binary scoring model.
-Appendix C The Dynamic Programming Approach used in KaSLA
-We illustrate the details of the dynamic programming algorithm used in KaSLA in Algorithm 1.
-Input: relevance score set ${\mathcal{R}}$, redundancy score set ${\mathcal{W}}$, the upper redundancy tolerance ${\mathcal{U}}$
-Output: selection set $\hat{\mathcal{H}}$
-$n\leftarrow|{\mathcal{R}}|$, $^{}^\hat{\mathcal{H}}\leftarrow\{\}$
-[…]
-[…]
-$keep\leftarrow\text{array of }(n+1)\times({\mathcal{U}}+1)$
-initialized to False
-for $i\leftarrow 1$ to $n$ do
-       for $w\leftarrow 0$ to ${\mathcal{U}}$ do
-             if ${\mathcal{W}}[i-1]\leq w$ then
-                   if $({\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]])>\mathcal{A}[i-1][w]$ then
-                         $\mathcal{A}[i][w]\leftarrow{\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]]$
-                         $keep[i][w]\leftarrow\text{True}$
-                  else
-                         $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
-            else
-                   $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
-$k\leftarrow{\mathcal{U}}$  for $i\leftarrow n$ downto $1$ do
-       if $keep[i][k]$ then
-             ^.$\hat{\mathcal{H}}.\text{add}(i-1)$
-             $k\leftarrow k-{\mathcal{W}}[i-1]$
-return $\hat{\mathcal{H}}$
-Algorithm 1 Dynamic Programming-based 0-1 Knapsack Problem Optimization
-Appendix D Complexity Analysis
-The efficiency of KaSLA stems from its hierarchical approach, dynamic programming and inherent parallelism.
-KaSLA’s time complexity is determined by two main phases: table linking and column linking within the linked tables.
-For a database with $n_{t}$ tables and a table capacity of $C_{t}$, the table linking phase has a time complexity of $O(n_{t}C_{t})$.
-KaSLA then links columns within the selected table.
-Let $n_{c}$ and $C_{c}$ represent the number of columns and the column capacity of a selected table, respectively.
-The time complexity of the column linking phase is $O(n_{c}C_{c})$.
-Consequently, the total time complexity of KaSLA is $O(n_{t}C_{t}+n_{t}(n_{c}C_{c}))$, indicating that the complexity of KaSLA scales relevant with the database dimensions.
-Appendix E The Ablation Studies for Different Language Models
-We present the ablation studies on the choice of language models for the probabilistic scoring model, as well as the performance scaling of various language models for the binary scoring model used in KaSLA.
-These studies are detailed in Table 8 and Table 7, respectively.
-Dataset BERT -Base BERT -Large RoBERTa -Base RoBERTa -Large SGPT -1.3B
-BIRD-dev 59.12 59.52 59.65 60.17 59.97
-Table 7: Execution Accuracy (EX) (%) of the scaling up performance about the choice of language model for binary scoring model.
-We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
-Dataset Deepseek-coder -1.3B Deepseek-coder -6.5B Deepseek-coder -7B
-BIRD-dev 60.17 61.28 61.47
-Table 8: Execution Accuracy (EX) (%) of the ablation study about the choice of language model for probabilistic scoring model.
-We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.